### v1

# Revisiting and Maximizing Temporal Knowledge in Semi-supervised Semantic Segmentation

Abstract
In semi-supervised semantic segmentation, the Mean Teacher- and co-training-based approaches are employed to mitigate confirmation bias and coupling problems. However, despite their high performance, these approaches frequently involve complex training pipelines and a substantial computational burden, limiting the scalability and compatibility of these methods. In this paper, we propose a PrevMatch framework that effectively mitigates the aforementioned limitations by maximizing the utilization of the temporal knowledge obtained during the training process. The PrevMatch framework relies on two core strategies: (1) we reconsider the use of temporal knowledge and thus directly utilize previous models obtained during training to generate additional pseudo-label guidance, referred to as previous guidance. (2) we design a highly randomized ensemble strategy to maximize the effectiveness of the previous guidance. Experimental results on four benchmark semantic segmentation datasets confirm that the proposed method consistently outperforms existing methods across various evaluation protocols. In particular, with DeepLabV3+ and ResNet-101 network settings, PrevMatch outperforms the existing state-of-the-art method, Diverse Co-training, by +1.6 mIoU on Pascal VOC with only 92 annotated images, while achieving 2.4 times faster training. Furthermore, the results indicate that PrevMatch induces stable optimization, particularly in benefiting classes that exhibit poor performance. Code is available at https://github.com/wooseok-shin/PrevMatch.
Index Terms:
Consistency regularization, semantic segmentation, semi-supervised learning, temporal knowledge
1 Introduction
Semantic segmentation [1, 2, 3, 4] is a critical task in various computer vision-related applications, such as autonomous driving [5, 6, 7], medical image analysis [8, 9], robotics [10], among others [11, 12, 13], where the goal is to assign a semantic class label to each pixel in an image. Despite the recent success of supervised learning-based methods in semantic segmentation, obtaining precise pixel-level annotations for supervised learning is extremely time-consuming and expensive. This limits the applicability of supervised learning-based methods across various domains and fields. Thus, extensive research has been conducted in semi-supervised semantic segmentation to overcome this limitation. The research has focused on developing methods enabling models to learn effectively from a limited number of labeled images along with a large number of unlabeled images.
In semi-supervised learning, self-training [14, 15, 16, 17, 18, 19, 20, 21] and consistency regularization [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] have become the predominant approaches for utilizing unlabeled data. In particular, self-training involves generating pseudo-labels using the predictions of the current model at each iteration for unlabeled samples and leveraging them to train the model in conjunction with labeled data. Consistency regularization encourages a network to predict consistently for various perturbed forms of identical input. Recent studies have focused on designing frameworks that combine self-training and consistency regularization to exploit the strengths of each method. However, self-training-based methods still suffer from the confirmation bias problem [17] even when combined with consistency regularization.
This problem is attributed to the accumulation of pseudo-label errors produced by the model itself, which exacerbate as self-training progresses. To mitigate this problem, existing methods distinguish the model prediction processes for supervised outputs and pseudo-labels. In other words, the supervised outputs and pseudo-labels are obtained from different predictions, respectively. Among them, some studies [34, 35, 32] have adopted the weak-to-strong consistency paradigm from the perspective of input separation, which supervises the prediction from a strongly perturbed input using the pseudo-label generated from its weakly perturbed counterpart (Fig. 1a). Its success is based on the idea that more reliable pseudo-labels can be derived from weak perturbations, and strong perturbations aid in mitigating confirmation bias [17, 21], broadening the knowledge and unlabeled data space [36, 32, 29], and shifting the model’s decision boundary to low-density regions [37]. Because of this advantage, the weak-to-strong consistency paradigm has become a fundamental component of the most recent methods.
Figure 1: Illustration of the frameworks for (a) FixMatch [34] (or PseudoSeg [35]), (b) Mean Teacher-based structure [26, 38, 39, 40], (c) Co-training [28, 29] (cps: cross pseudo supervision), (d) Dual Mean Teacher [30, 31], and (e) the proposed method. In (d), the inputs ($x^{s_{cut}},x^{s_{class}}$) indicate the CutMix [41] and ClassMix [42] augmentations used in Dual Teacher [31].
Another method to obtain different prediction views involves using network perturbation based on a teacher-student structure. In particular, Mean Teacher [26] is a representative approach in semi-supervised segmentation, where a teacher network is derived using an exponential moving average (EMA) of the student model’s weights (Fig. 1b). Although the Mean Teacher generates somewhat different prediction views between the teacher and student, this method is limited by a coupling problem [27]: as training progresses, the teacher and student become tightly linked, and consequently, the teacher’s predictions become similar to those of the student. To mitigate the coupling problem and obtain diverse predictions, certain studies [30, 31] have proposed a dual EMA teacher-based framework where two teachers are alternatively updated at every epoch (Fig. 1d). Among them, PS-MT [30] produces more reliable pseudo-labels by ensembling the predictions of the two teachers. By contrast, Na et al. [31] reported that ensembling predictions can reduce the diversity of pseudo-labels. Thus, they proposed a Dual Teacher framework that alternately activates two teachers in each epoch to generate diversified pseudo-labels. Although these studies have demonstrated the benefits of using multiple teachers to mitigate the coupling problem, their ability to provide reliable and diverse pseudo-labels simultaneously remains limited. In addition, they have incorporated additional complex components to ensure diversity between the two teachers. In particular, adversarial feature perturbation and a new loss function are used for PS-MT. Further, distinct types of augmentation provided to each teacher (e.g., CutMix or ClassMix) and layer perturbation are used for Dual Teacher. Consequently, this complexity can hinder the scalability and compatibility of these approaches with existing semi-supervised methods.
Instead of using EMA-based teachers, a co-training paradigm [43, 44, 45, 46, 27, 28, 47, 29] has been widely used to expand prediction views (Fig. 1c). This paradigm involves simultaneously training multiple networks with different initializations in a mutual teaching manner, where each network supervises the others using pseudo-labels generated from its predictions. Building on this concept, subsequent studies [45, 29] have demonstrated that increasing the diversity of pseudo-label views from student networks improves their generalization ability. For example, this can involve the use of more co-networks, diverse input domains (e.g., RGB and frequency), or different architectures (e.g., CNN and Transformer). Co-training provides diverse pseudo-label guidance with stability and without concerns regarding the coupling problem. However, its scalability remains constrained due to computational complexity and resource demands.
In this paper, we propose the PrevMatch framework, which efficiently expands pseudo-label views by maximizing the utilization of previous models obtained during training, as depicted in Fig. 1e. The PrevMatch framework is based on two main ideas. First, to efficiently address the coupling problem, we revisit the utilization of temporal knowledge. Specifically, we save several models at specific epochs during training and utilize their predictions as additional guidance, referred to as previous guidance, which acts as a regularizer in conjunction with standard guidance. This strategy addresses the coupling problem and reduces the complexity associated with additional training components. Second, we design a highly randomized ensemble strategy to maximize the effectiveness of utilizing the previous guidance. This approach involves selecting a random number of models from those previously saved and ensembling their predictions using randomized weights. This strategy can efficiently provide diverse and reliable pseudo-labels, while avoiding the significant computational complexity inherent in co-training approaches.
Extensive experiments conducted across various evaluation protocols on the PASCAL, Cityscapes, COCO, and ADE20K datasets reveal that the proposed PrevMatch significantly outperforms existing methods. In particular, compared to Diverse Co-training, the current state-of-the-art method, PrevMatch achieves a +1.6 mIoU improvement on Pascal VOC with 92 labels while accelerating training by 2.4 times. In addition to quantitative evaluations, ablation studies and analyses of the proposed components are conducted to explain the success of PrevMatch.
2 Related Work
2.1 Semi-supervised Learning
The central challenge in semi-supervised learning lies in extracting additional supervision for unlabeled samples. Previous studies can be categorized into two main approaches: self-training and consistency regularization. Self-training-based methods [14, 15, 16, 17, 18, 19, 20] assign pseudo-labels to unlabeled data using the ongoing model predictions and integrate them with labeled data to retrain the model. Consistency regularization methods [22, 23, 24, 25, 34, 48, 26, 43, 27, 49] enforce a model to produce invariant predictions for different perturbations, such as input and network perturbations.
In terms of input perturbations [22, 23, 24, 25, 34, 48], the $\Pi$-model [50, 22, 23] applies simple stochastic perturbations, while VAT [24] applies adversarial perturbations to unlabeled samples. Subsequently, UDA [25] expands the perturbation pool using RandAugment [51] to provide diverse augmented samples. Furthermore, FixMatch [34], a standard approach in recent semi-supervised learning, adopts a weak-to-strong paradigm where the prediction for a strongly perturbed input is supervised by pseudo-labels generated from the prediction for a weakly perturbed input. Network perturbation, another approach to enforcing consistency regularization, has also been widely investigated [23, 26, 43, 27, 49]. Temporal Ensembling [23] and Mean Teacher [26] methods leverage earlier predictions and weights of the model, respectively, to obtain different prediction views. In addition, co-training-based methods [43, 27, 49] adopt multiple networks with different initializations and enforce consistency regularization between their predictions. Among the numerous studies discussed, the weak-to-strong paradigm, Mean Teacher structure, and co-training approaches have been widely used to advance semi-supervised semantic segmentation.
2.2 Semi-supervised Semantic Segmentation
Owing to the detailed nature of the segmentation task, which involves numerous pixels and multiple classes within an image, semantic segmentation requires more time-consuming and costly efforts to annotate labels and demands sophisticated approaches for accurate performance. Thus, recent research has increasingly focused on developing methods for semi-supervised semantic segmentation, based on previous achievements in semi-supervised learning. Early studies [52, 53] employed a GAN framework [54] to generate additional supervision for unlabeled images by distinguishing between pseudo-labels and manual labels. Recent studies have focused on developing improved methods based on weak-to-strong consistency, Mean Teacher, and co-training paradigms. PseudoSeg [35] adopts a weak-to-strong paradigm based on a single network. CPS [28, 45] and GCT [55] employ two networks for co-training and demonstrate that co-training outperforms the Mean Teacher approach. Subsequent studies have integrated the weak-to-strong paradigm (i.e., input perturbation) into the Mean Teacher or co-training approaches (i.e., network perturbation), demonstrating performance improvements.
A stream of research that combines the weak-to-strong paradigm with the Mean Teacher structure has proposed techniques such as advanced data augmentation [38, 56, 39, 40], prototype learning [57], curriculum learning [58], symbolic reasoning [59], and dual Mean Teacher [30, 31]. Among them, PS-MT [30] and Dual Teacher [31] methods implement a dual EMA teacher-based framework to mitigate the coupling problem between the teacher and student models, with two teachers alternately updating each epoch based on the EMA of the student’s weights. However, their pipelines inevitably involve complex components to resolve the coupling problem, limiting the scalability and compatibility of these approaches. Moreover, regarding pseudo-label generation, PS-MT ensembles the predictions of two teachers to improve reliability, whereas Dual Teacher alternately uses the predictions of each teacher to enhance diversity. However, neither method satisfies both reliability and diversity. In other words, PS-MT improves the reliability of pseudo-labels but lacks diversity, whereas Dual Teacher suffers from the opposite limitation.
Another stream of research, combining the weak-to-strong paradigm with co-training, has developed techniques such as a shared backbone with multiple heads [46], conservative-progressive learning [60], and increasing the diversity of co-training [29]. Among them, Li et al. [29] investigated the working mechanism of co-training and discovered that providing distinct pseudo-label views improves generalization ability. To this end, they proposed Diverse Co-training, which incorporates variations in the input domains (RGB and frequency) and architectures (CNN and Transformer). However, in diverse co-training, the number of whole networks to train increases as pseudo-label views increase.
In this study, instead of using Mean Teacher or co-training approaches, we revisit the utilization of temporal knowledge and efficiently expand pseudo-label views by maximizing the use of previous models. The proposed PrevMatch framework can be seen as simplifying and extending [30, 31] while also enhancing the efficiency of [29]. Specifically, we eliminate the complex components used in [30, 31], such as distinct augmentation types for each epoch, layer/adversarial feature perturbations, EMA teachers, and a new loss function. Instead, PrevMatch reuses the previous models and the weakly perturbed input used in the standard flow, thereby improving the simplicity and compatibility of the overall framework. Furthermore, we provide reliable and diverse pseudo-labels to the student network through a highly randomized ensemble strategy. Moreover, although both PrevMatch and co-training provide diverse pseudo-label views, PrevMatch operates on a single trainable network with fixed previous models. This facilitates greater efficiency in terms of computational and memory costs.
2.3 Temporal Knowledge in Semi-supervised Learning
In the context of semi-supervised learning, several studies have leveraged temporal knowledge obtained from previous training stages. Temporal Ensembling [23] accumulates predictions for unlabeled samples across different epochs using the EMA approach and enforces consistency between the current and EMA predictions. Mean Teacher [26] averages the model weights across the training steps using the EMA approach, producing a more stable teacher network. As a further extension, a dual EMA teacher-based framework was proposed [30, 31], where two teachers are alternately updated at each epoch. Moreover, TC-SSL [20] measures the time-consistency (TC) scores of individual samples across training epochs and selects unlabeled samples with higher TC scores for consistency learning. This approach assumes that time-consistent predictions are typically accurate. Similar to TC-SSL, ST++ [21] proposes a selective re-training scheme that selects more reliable samples based on a stability score. The stability score based on the mIoU metric is derived from comparisons between the predictions of the final model and those of several checkpoints obtained during training. In addition to semi-supervised learning, Feng et al. [61] proposed a temporal consistency framework that adopts two temporary teachers to learn instance temporal consistency in representation learning.
3 Method
This section describes the preliminaries and overall training flow for semi-supervised semantic segmentation. The PrevMatch method is also introduced here.
3.1 Preliminaries & Overall Workflow
Semi-supervised semantic segmentation aims to fully utilize unlabeled images $D_{u}=\{x^{u}_{i}\}$, given only a limited number of labeled images $D_{l}=\{(x^{l}_{i},y^{l}_{i})\}$. In general semi-supervised learning, the objective function is divided into supervised loss $L_{s}$ and unsupervised loss $L_{u}$ as follows:
$L=\frac{1}{2}(L_{s}+L_{u}).$ (1)
Figure 2: Overall framework of PrevMatch (based on UniMatch). FP indicates the feature perturbation used in UniMatch. $g$ and $h$ are the encoder and decoder of the entire network $f$, respectively. When the maximum length of the previous list is exceeded, the oldest model is replaced with the new model using a first-in-first-out (FIFO) approach. The red and blue arrows denote standard and proposed guidance flows, respectively.
  1. 1.
    Supervised Flow In supervised flow, a segmentation network $f$ receives a labeled image $x^{l}$ and generates the corresponding predicted class distribution $p^{l}$. Subsequently, the supervised loss is calculated using pixel-wise cross-entropy $H$ between the prediction and the ground-truth label. This can be formulated as follows.
    $L_{s}=\sum_{i,j}H(y^{l}_{ij},p^{l}_{ij}),$ (2)
    where $i$ and $j$ indicate pixel indices.
  2. 2.
    Unsupervised – Standard Flow Following the success of the weak-to-strong consistency paradigm popularized by FixMatch [34], most semi-supervised segmentation methods have adopted this paradigm. Concretely, two perturbed images (i.e., $x^{w}$ and $x^{s}$) are obtained by applying weak and strong augmentations to an unlabeled image $x^{u}$. The network $f$ receives the two perturbed images and outputs the predicted class distributions $p^{w}$ and $p^{s}$. As there is no ground-truth label for the unlabeled image, a pseudo-label for the weakly perturbed prediction is obtained by: $y^{w}\text{= argmax}(p^{w})$. Then, this pseudo-label is used to supervise the strongly perturbed prediction, as depicted in the standard flow presented in Fig. 1e (red arrow). This consistency term can be formulated as follows.
    $C(p^{w},p^{s})=\sum_{i,j}\mathbb{1}(max(p^{w}_{ij})\geq\tau)H(y^{w}_{ij},p^{s}_{ij}),$ (3)
    where $\mathbb{1}(\cdot)$ is an indicator function, and $\tau$ is a confidence threshold used to ignore noise in a pseudo-label. Furthermore, the UniMatch framework [32] introduced two contributions based on FixMatch: the dual stream image-level perturbation and the feature-level perturbation strategies, to expand the perturbation space, thereby achieving significant improvements. Therefore, we select the UniMatch framework as our baseline. In particular, two strongly perturbed images ($x^{s_{1}}$ and $x^{s_{2}}$) are derived by randomly applying a strong augmentation pool to the same image, $x^{u}$. Subsequently, their corresponding predictions ($p^{s_{1}}$ and $p^{s_{2}}$) are generated through the network $f$. Regarding the feature-level perturbation strategy, the prediction is obtained as follows: $p^{fp}=h(\text{Dropout}(g(x^{w})))$, where $g$ and $h$ are the encoder and decoder of the entire network $f$, respectively. Ultimately, three predictions are simultaneously supervised by a common pseudo-label derived from a weak view (red arrow in Fig. 2). This can be formulated as follows.
    $L_{u(standard)}=C(p^{w},p^{s_{1}})+C(p^{w},p^{s_{2}})+C(p^{w},p^{fp}).$ (4)
  3. 3.
    Unsupervised – Proposed Flow To obtain additional pseudo-label guidance, the same weakly perturbed image, $x^{w}$, used in the standard branch, is fed into the previous model branch. Subsequently, $k$ models, where $k$ is chosen randomly from $\{1,2,...,K\}$, are selected from the saved list of previous models, and $k$ predictions for $x^{w}$ are generated. Previous guidance, $p^{w^{\prime}}$, is obtained by aggregating the $k$ predictions using randomized ensemble weights to improve the diversity of the pseudo-label view. As illustrated by the blue arrow in Fig. 2, the previous guidance functions as additional regularizers, supervising the three predictions. This can be formulated as follows.
    $L_{u(prev)}=C(p^{w^{\prime}},p^{s_{1}})+C(p^{w^{\prime}},p^{s_{2}})+C(p^{w^{\prime}},p^{fp}).$ (5)
Finally, the total loss of the unsupervised flow is defined by combining the standard and proposed flow losses as follows:
$L_{u}=L_{u(standard)}+\lambda\cdot L_{u(prev),}$ (6)
where $\lambda$ denotes the weight of the proposed flow.
3.2 Previous Guidance: Revisiting Temporal Knowledge
To address the coupling problem between student and teacher networks, we revisit the utilization of temporal knowledge that can be obtained during the training process. In the literature, Temporal Ensembling [23] and Mean Teacher [26] methods average the temporal knowledge of a model in terms of its predictions and weights, respectively. PS-MT [30] and Dual Teacher [31], inheriting the spirit of the Mean Teacher [26] method, aim to further exploit temporal knowledge by adopting two EMA teachers. TC-SSL [20] and ST++ [21] methods implement a filtering mechanism to exclude less-informative unlabeled samples using specific scoring criteria measuring temporal consistency.
In contrast to existing methods, we directly utilize previous models and generate pseudo-label guidance from their predictions without relying on a complex pipeline of dual EMA-based methods or a filtering mechanism. In particular, we store multiple models at different epochs that meet the specified criteria during training and use their pseudo-labels as additional guidance, referred to as previous guidance. As training progresses, the decoupling between the student and previously saved teachers increases, allowing for the acquisition of different prediction views. However, as decoupling becomes more pronounced, the positive effects of self-training from correct pseudo-labels may diminish due to the use of outdated teachers. Therefore, previous guidance is used in conjunction with the standard guidance obtained from the current student. In addition, we define the maximum length ($N$) of the list for storing previous models and replace the oldest teacher with a newer one when this limit is exceeded to avoid using excessively outdated teachers. Formally, one previous model is randomly selected from the previous model list $\{T_{1},T_{2},...,T_{N}\}$. Then, this model processes the same weakly perturbed image, $x^{w}$, used in the standard flow and produces predictions $p^{w^{\prime}}$. The previous guidance is obtained by: $y^{w^{\prime}}\text{= argmax}(p^{w^{\prime}})$, and it supervises three predictions according to Eq. (5). In this way, we produce diverse pseudo-labels, as in prior studies [30, 31, 28, 29], by leveraging different previous models encompassing varied perspectives of temporal knowledge, without complex additional components or heavy computational burden.
Save Criteria. In this approach, storing the appropriate previous models is crucial for generating diverse and reliable pseudo-labels. Some studies [62, 63, 21] that employ intermediate models for ensembling or filtering in image recognition and segmentation save the model at regular intervals (e.g., every 20 epochs in a total of 100 epochs). By contrast, we save the model when it achieves the best performance on the validation set to ensure the stability of the previous guidance. The weights and performances of the neural network can vary significantly during the optimization process, and excessively large fluctuations may increase negative impacts. These negative impacts can be exacerbated by label scarcity and a pronounced class imbalance in semi-supervised semantic segmentation compared to standard image recognition. In addition, the periodic saving method requires additional hyperparameter searches to determine appropriate intervals. Therefore, we adopt our saving approach to achieve stability and simplicity (refer to Table IX for related experiments).
The efficacy of the previous guidance can be intuitively explained as follows. When using only standard guidance, two training scenarios arise based on whether the prediction is correct or incorrect. In cases where the prediction is incorrect, the network is trained in the wrong direction. In contrast, given the standard and previous guidance, four scenarios can be considered based on whether they are correct or incorrect (standard-previous): (1) correct-correct, (2) correct-incorrect, (3) incorrect-correct, and (4) incorrect-incorrect. Through case (3), the network receives an additional opportunity to be guided in the right direction, away from the wrong one. Although model training may be hindered through case (2), leading to significant fluctuations, we empirically demonstrate that the positive effects of the proposed method outweigh the negative effects (which is further supported by the analysis of training stability in Section 4.4.1). In addition, previous guidance can help mitigate the network’s catastrophic forgetting problem [64, 65], where previously learned knowledge is forgotten when acquiring new knowledge. In particular, this phenomenon can be more pronounced in self-training and class-imbalance scenarios (i.e., in our scenario) due to the lack of labeled data, potentially leading to significant performance fluctuations in poorly behaved classes [19]. In semi-supervised [20] and representation learning [61], some studies have demonstrated that utilizing temporal knowledge helps mitigate the catastrophic forgetting problem and stabilizes training. Based on this fact, we explore the learning stability of the proposed method in Section 4.4.1.
3.3 Maximizing Efficacy of Previous Guidance
One way to improve the reliability of the predictions involves using network ensemble techniques. These techniques have been widely used in various domains as a promising method for improving performance. In particular, several studies [62, 63] in the field of image recognition have demonstrated that ensembles of multiple intermediate models obtained during training also improve prediction accuracy and diversity. Therefore, we utilize network ensembling to improve the reliability of the previous guidance. In designing this method, we also consider computational complexity and pseudo-label diversity. To this end, given the list that includes $N$ previous models, we randomly select $K$ ($K\leq N$) models for each iteration to mitigate the increase in computational complexity while ensuring pseudo-label diversity. However, this approach, which always ensembles $K$ models, may not guarantee the diversity of pseudo-labels, as noted in Dual Teacher [31]. This problem can worsen for large values of $K$.
Therefore, to provide reliable and diverse pseudo-labels to the student network, we propose a highly randomized ensemble strategy comprising the following two ideas:
  • Random Selection: For each iteration, we randomly select a varying number of teachers, $k$, ranging from 1 to $K$. For example, $k$=1,2, or 3 can be selected for each iteration when $K$=3. With this approach, selecting a large $k$ tends to yield consistent pseudo-labels, enhancing reliability, while a small $k$ contributes to increased diversity in the pseudo-labels. In addition, computational costs are lower than those incurred when using a fixed number $K$. This strategy enables the student network to obtain stable and diverse guidance, which functions as a robust regularizer, aiding network optimization.
  • Random Weights: Regarding ensemble weights, we propose a random aggregation strategy that averages $k$ predictions using random weights for each iteration instead of a simple average of $k$ predictions. In particular, the $k$ selected teachers receive $x^{w}$ and output $k$ predictions, $\{p^{w^{\prime}}_{1},p^{w^{\prime}}_{2},...,p^{w^{\prime}}_{k}\}$. The final previous guidance is then obtained by aggregating these predictions using random weights as follows:
    $p^{w^{\prime}}=\sum_{i=1}^{k}w_{i}\cdot p^{w^{\prime}}_{i},$ (7)
    where $w_{i}$ is derived from a Dirichlet distribution as follows: $\{w_{1},w_{2},...,w_{k}\}\sim\text{Dir}(\alpha_{1},\alpha_{2},...,\alpha_{k})$. Note that the sum of $w_{i}$ is one. This approach explores all combinations of previous guidance in a continuous space, expanding the original pseudo-label space beyond a simple average.
In this way, the proposed randomized ensemble strategy can improve the reliability and diversity of the previous guidance while mitigating the increase in computational complexity.
4 Experiments
4.1 Experimental Configuration
4.1.1 Datasets
PASCAL VOC [66] is a widely used benchmark dataset in semantic segmentation, comprising 21 distinct categories (20 object types and a background class). This dataset contains 10,582 images designated for training purposes, segmented into two subsets based on annotation quality: 1,464 images constitute the high-quality subset, characterized by detailed annotations, whereas the remaining 9,118 images form the coarse subset, featuring less detailed annotations. This study follows established protocols for a fair comparison [28, 67, 32, 39]. In particular, three training protocols are considered based on the criteria for selecting labeled images.
  • Original: A protocol whereby labeled images are exclusively sourced from the high-quality subset.
  • Blended: A protocol that entails a random selection of labeled images from the total dataset.
  • Priority: A protocol where the selection of labeled images is first derived from the high-quality subset; if not sufficient, it is complemented by additional images from the coarse subset.
Cityscapes dataset [68], tailored for the semantic analysis of urban street scenes, comprises 2,975 high-resolution images for training and 500 images for validation, primarily focusing on 19 categories within urban environments. Consistent with previous studies [21, 32, 29], we evaluate this dataset using various label partitions, specifically 1/16, 1/8, 1/4, and 1/2 of the total number of labels.
COCO dataset [69], notable for its complexity and scale, comprises 118k training and 5k validation images. This dataset features dense annotations across 81 classes set in various indoor and outdoor scenarios. Considering the performance plateaus observed in datasets such as Pascal and Cityscapes and the higher number of classes, the COCO dataset has emerged as a practical and valuable benchmark for evaluating advanced algorithms in semi-supervised segmentation. Following existing studies [32], we validate the proposed method using 1/256, 1/128, and 1/64 label partitions.
ADE20K dataset [70], containing more diverse scenes and a greater number of classes (150 categories) than COCO, comprises 20,210 images for training and 2,000 images for validation. Using the label partitions from existing studies [33], we evaluate the proposed method on 1/128, 1/64, and 1/32 partitions.
In all protocols, training images not selected as labeled images are utilized as unlabeled images.
4.1.2 Architecture and Implementation Details
Consistent with existing literature, we employ ResNet-50 and ResNet-101 backbones [71] as encoders for the Pascal VOC, Cityscapes, and ADE20K datasets. For the COCO dataset, Xception-65 [72] is adopted as the backbone. We use DeepLabV3+ [3] as the segmentation head and set its output stride to 16 for efficient training.
For the training setups, each mini-batch comprises 8 labeled and 8 unlabeled images. The proposed method is trained for 80, 240, 30, and 40 epochs for the PASCAL, Cityscapes, COCO, and ADE20K datasets, respectively, using the SGD optimizer. The learning rates are initially set to 0.001, 0.005, 0.004, and 0.004 for these datasets, respectively, and are managed using a polynomial learning rate scheduler. Moreover, we use 321/513, 801, 513, and 513 random crops for these datasets, respectively. For image augmentation, we use common weak (e.g., resize, crop, and flip) and strong (e.g., color transformations, grayscale, cutmix, and blur) data augmentations, as in UniMatch [32]. For the hyperparameters of standard consistency flow, $\tau$ is set to 0 for Cityscapes and 0.95 for the other datasets. The dropout rate for feature perturbation is set to 0.5.
For the hyperparameters of PrevMatch, $\tau$ is set to 0.9 for Pascal, 0 for Cityscapes, and 0.95 for COCO and ADE20k. The maximum length ($N$) of the previous list is set to eight for Pascal and Cityscapes and five for COCO and ADE20K. The upper bound number $K$ for the random selection is set to three. For weight $\lambda$, as previous guidance should correct the model before it becomes overfitted in the wrong direction (i.e., confirmation bias [17]) during the middle of training, setting an appropriate weight $\lambda$ is crucial. Additionally, as the initial model typically exhibits poor performance, we implement a warmup schedule for the weight of the proposed flow, similar to the commonly used polynomial learning rate scheduler that includes a warmup phase. Experiments are conducted in the following environments: UBUNTU 20.04, Python 3.10.4, PyTorch 1.12.1, CUDA 11.3, and NVIDIA RTX 3090Ti or A6000 GPUs. The training is performed using one GPU for Pascal and two GPUs for the other datasets.
TABLE I: Comparison with state-of-the-art methods on the Original protocol of Pascal VOC dataset. The numeric values in the header (e.g., 92) represent the number of labeled images used for training. All methods are trained using ResNet-50/101 and DeepLabV3+. The number of trainable whole networks and input resolution are reported. The values of PrevMatch are averaged over three runs. The evaluation metric is the mean IoU (%).
Pascal [Original set] Encoder #Trainable Networks Resolution # Labeled images (Total: 10582)
92 183 366 732 1464
Supervised Baseline R-50 $\times$1 5132 44.0 52.3 61.7 66.7 72.9
PseudoSeg [35] [ICLR’21] R-50 $\times$1 5132 54.9 61.9 64.9 70.4 71.0
PC2Seg [73] [ICCV’21] R-50 $\times$1 5132 56.9 64.6 67.6 70.9 72.3
CPCL [60] [TIP’23] R-50 $\times$2 5122 61.9 67.0 72.1 74.3 -
AugSeg [39] [CVPR’23] R-50 $\times$1 5122 64.2 72.2 76.2 77.4 78.8
UniMatch [32] [CVPR’23] R-50 $\times$1 3212 71.9 72.5 76.0 77.4 78.7
Dual Teacher [31] [NeurIPS’23] R-50 $\times$1 3212 70.8 74.5 76.4 77.7 78.2
PrevMatch (ours) R-50 $\times$1 3212 73.4 75.4 77.5 78.6 79.3
Supervised Baseline R-101 $\times$1 5132 45.1 55.3 64.8 69.7 73.5
CPS [28] [CVPR’21] R-101 $\times$2 5122 64.1 67.4 71.7 75.9 -
ReCo [74] [ICLR’22] R-101 $\times$1 3212 64.8 72.0 73.1 74.7 -
PS-MT [30] [CVPR’22] R-101 $\times$1 5122 65.8 69.6 76.6 78.4 80.0
ST++ [21] [CVPR’22] R-101 $\times$1 3212 65.2 71.0 74.6 77.3 79.1
U2PL [67] [CVPR’22] R-101 $\times$1 5122 68.0 69.2 73.7 76.2 79.5
GTA-Seg [75] [NeurIPS’22] R-101 $\times$2 5132 70.0 73.2 75.6 78.4 80.5
PCR [57] [NeurIPS’22] R-101 $\times$1 5132 70.1 74.7 77.2 78.5 80.7
DGCL [76] [CVPR’23] R-101 $\times$1 5132 70.5 77.1 78.7 79.2 81.6
CCVC [47] [CVPR’23] R-101 $\times$2 5122 70.2 74.4 77.4 79.1 80.5
iMAS [40] [CVPR’23] R-101 $\times$1 5132 68.8 74.4 78.5 79.5 81.2
AugSeg [39] [CVPR’23] R-101 $\times$1 5122 71.1 75.5 78.8 80.3 81.4
UniMatch [32] [CVPR’23] R-101 $\times$1 3212 75.2 77.2 78.8 79.9 81.2
ESL [58] [ICCV’23] R-101 $\times$1 5132 71.0 74.1 78.1 79.5 81.8
LogicDiag [59] [ICCV’23] R-101 $\times$1 5132 73.3 76.7 77.9 79.4 -
LogicDiag + MKD [59] [ICCV’23] R-101 $\times$2 5132 74.7 77.2 78.4 80.1 -
Diverse Co-T. (2-cps) [29] [ICCV’23] R-101 $\times$2 3212 74.8 77.6 79.5 80.3 81.7
Diverse Co-T. (3-cps) [29] [ICCV’23] R-101 $\times$3 3212 75.4 76.8 79.6 80.4 81.6
PrevMatch (ours) R-101 $\times$1 3212 77.0 78.5 79.6 80.4 81.6
TABLE II: Comparison with state-of-the-art methods on the Blended protocol of Pascal VOC dataset. All methods are trained using ResNet-50 and DeepLabV3+. The fractional values indicate the ratio of labeled images used for training.
Pascal [Blended set] Resolution 1/16 1/8 1/4
Supervised Baseline 5132 62.4 68.2 72.3
Mean Teacher [26] [NeurIPS’17] 5122 66.8 70.8 73.2
CCT [44] [CVPR’20] 5122 65.2 70.9 73.4
GCT [55] [ECCV’20] 5122 64.1 70.5 73.5
CutMix-Seg [38] [BMVC’20] 5122 68.9 70.7 72.5
CAC [77] [CVPR’21] 3202 70.1 72.4 74.0
CPS [28] [CVPR’21] 5122 72.0 73.7 74.9
UCC [46] [CVPR’22] 5122 74.1 74.8 76.4
PS-MT [30] [CVPR’22] 5122 72.8 75.7 76.4
UniMatch [32] [CVPR’23] 3212 74.5 75.8 76.1
AugSeg [39] [CVPR’23] 5122 74.7 76.0 77.2
iMAS [40] [CVPR’23] 5132 74.8 76.5 77.0
CCVC [47] [CVPR’23] 5122 74.5 76.1 76.4
PrevMatch (ours) 3212 75.6 76.4 76.3
5132 76.0 77.1 77.6
TABLE III: Comparison with state-of-the-art methods on the Priority protocol of Pascal VOC dataset. All methods are trained using ResNet-101 and DeepLabV3+.
Pascal [Priority set] Resolution 1/16 1/8 1/4
Supervised Baseline 5132 70.6 75.0 76.5
U2PL [67] [CVPR’22] 5122 77.2 79.0 79.3
UniMatch [32] [CVPR’23] 5132 80.9 81.9 80.4
AugSeg [39] [CVPR’23] 5122 79.3 81.5 80.5
Dual Teacher [31] [NeurIPS’23] 5132 80.1 81.5 80.5
PrevMatch (ours) 5132 81.4 81.9 80.8
4.2 Comparison with State-of-the-Art Methods
PASCAL VOC.   Three main experiments are conducted on the Original, Blended, and Priority protocols, which are divided according to the criteria for selecting labeled images. In Table I, the proposed method outperforms state-of-the-art methods in almost all partitions, even with a single trainable network and a lower resolution. In particular, PrevMatch significantly improves performance in settings with fewer labels (92 and 183). Compared to Diverse Co-T, PrevMatch exhibits comparable performance in settings with more than 366 labels, but it operates on a single trainable network, indicating that the proposed method is efficient with respect to training costs. For Blended and Priority protocols reported in Tables II and III, the proposed method consistently outperforms existing methods.
Cityscapes.   Experimental results for the Cityscapes validation set are listed in Table IV. Under four label partitions using ResNet-50 and ResNet-101 backbones, the proposed method outperforms previous methods in six out of eight cases. Similar to the results reported in Table 1, PrevMatch consistently improves performance in setups with fewer labels. However, certain methods (e.g., AugSeg, AEL, and Dual Teacher) perform better than PrevMatch at the 1/2 label setups. This can be attributed to the advanced data augmentation techniques used in these methods. In other words, these methods could achieve better performance because the 1/2 label partition increasingly resembles a supervised learning setting, where augmentation techniques play a crucial role. Thus, we intend to explore the efficacy of the advanced augmentations on the proposed method in future work.
TABLE IV: Comparison with state-of-the-art methods on the Cityscapes dataset. All methods are trained using ResNet-50/101 and DeepLabV3+.
Cityscapes Encoder #Trainable 1/16 1/8 1/4 1/2
Networks (186) (372) (744) (1488)
Supervised Baseline R-50 $\times$1 63.3 70.2 73.1 76.6
PS-MT [30] [CVPR’22] R-50 $\times$1 - 75.8 76.9 77.6
U2PL [67] [CVPR’22] R-50 $\times$1 70.6 73.0 76.3 77.2
UniMatch [32] [CVPR’23] R-50 $\times$1 75.0 76.8 77.5 78.6
iMAS [40] [CVPR’23] R-50 $\times$1 74.3 77.4 78.1 79.3
AugSeg [39] [CVPR’23] R-50 $\times$1 73.7 76.5 78.8 79.3
CCVC [47] [CVPR’23] R-50 $\times$2 74.9 76.4 77.3 -
FPL (w/ CPS) [78] [CVPR’23] R-50 $\times$2 74.8 77.3 78.5 -
Diverse Co-T. (3-cps) [29] [ICCV’23] R-50 $\times$3 - 76.5 77.9 -
PrevMatch (ours) R-50 $\times$1 75.8 77.8 78.8 79.2
Supervised Baseline R-101 $\times$1 66.3 72.8 75.0 78.0
CPS [28] [CVPR’21] R-101 $\times$2 69.8 74.3 74.6 76.8
AEL [56] [NeurIPS’21] R-101 $\times$1 75.8 77.9 79.0 80.3
PS-MT [30] [CVPR’22] R-101 $\times$1 - 76.9 77.6 79.1
U2PL [67] [CVPR’22] R-101 $\times$1 74.9 76.5 78.5 79.1
PCR [57] [NeurIPS’22] R-101 $\times$1 73.4 76.3 78.4 79.1
CISC-R [79] [TPAMI’23] R-101 $\times$1 - 75.9 77.7 -
FPL (w/ AEL) [78] [CVPR’23] R-101 $\times$1 76.6 78.2 78.5 -
AugSeg [39] [CVPR’23] R-101 $\times$1 75.2 77.8 79.6 80.4
UniMatch [32] [CVPR’23] R-101 $\times$1 76.6 77.9 79.2 79.5
ESL [58] [ICCV’23] R-101 $\times$1 75.1 77.2 78.9 80.5
UPC (w/ U2PL) [80] [ICCV’23] R-101 $\times$1 75.3 77.4 79.0 79.6
Diverse Co-T. (2-cps) [29] [ICCV’23] R-101 $\times$2 75.0 77.3 78.7 -
Diverse Co-T. (3-cps) [29] [ICCV’23] R-101 $\times$3 75.7 77.4 78.5 -
Dual Teacher [31] [NeurIPS’23] R-101 $\times$1 76.8 78.4 79.5 80.5
PrevMatch (ours) R-101 $\times$1 77.7 78.9 80.1 80.1
COCO & ADE20K.   Table V lists results for the large-scale datasets COCO and ADE20K. The proposed method consistently improves the performance across all partitions compared to the baseline, UniMatch, suggesting that PrevMatch is also effective on large-scale datasets.
TABLE V: Evaluation results on the large-scale datasets using Xception-65 for COCO and ResNet-50 for ADE20K.
Method COCO ADE20K
1/256 1/128 1/64 1/128 1/64 1/32
Supervised 28.0 33.6 37.8 7.2 9.9 13.7
UniMatch [32] 38.9 44.4 48.2 13.6 18.3 23.9
PrevMatch (ours) 40.2 45.7 48.4 15.4 19.6 24.9
4.3 Ablation Studies
In the ablation studies and discussion, the baseline refers to the UniMatch method [32].
4.3.1 Individual Efficacy of the Proposed Components
The effects of the individual components are investigated, and the results are presented in Table VI. The first row indicates the baseline. All components of the proposed method consistently achieve performance gains. In particular, previous guidance, which randomly selects one previous model from the previous list for each iteration to generate additional guidance, surpasses the baseline by 0.8% and 1.3% in the 92- and 183-label settings, respectively. Regarding the number of models ($K$) for the network ensemble, we experiment with simple ensemble (fixed $K$) or random selection (random $K$) strategies. The result (third row) obtained using fixed $K$ exhibits minor performance gains. In contrast, the results using random $K$ (fourth row) indicate significantly improved performance. This implies that a fixed $K$ ensemble improves the reliability of pseudo-labels but limits the diversity, as noted in Dual Teacher [31], whereas a random $K$ strategy can provide reliable and diverse guidance to the model. In addition, using random weights for network ensembling contributes to performance gains, indicating that it expands the original pseudo-label space by generating diversified guidance.
TABLE VI: Ablation study of the components of the proposed method using a ResNet-50 encoder. For the ensemble, we set $K$=3. The means and standard deviation are also reported based on three runs to validate statistical significance.
Previous Simple Random Random PASCAL
Guidance Ensemble Selection Weights 92 183
- - - - 71.9±0.5 72.5±0.7
- - - 72.7±0.5 73.8±0.6
- - 72.7±0.4 74.1±0.5
- - 73.2±0.4 74.9±0.6
- 73.4±0.4 75.4±0.5
4.3.2 Previous List Length
We investigate the effect of the length ($N$) of the previous list that stores the temporal models. In Table VII, the results for $N=1$ and $N=2$ are comparable to those of the baseline. This suggests that the aforementioned coupling problem may persist because the previous models in the list are continually updated with the latest model when $N$ is small. The cases of $N=$ 4, 8, and 12 consistently outperform the baseline, revealing that the proposed method is not highly sensitive to hyperparameters. However, the performance for $N=20$ increases only marginally due to the use of outdated teachers. Based on this result, we recommend setting the value of $N$ to approximately 5–15% of the total training epochs, which proves to be appropriate for different datasets (e.g., Pascal=6–10, Cityscapes=8–16, and COCO and ADE20k=4–5).
TABLE VII: Ablation study for the maximum length of the previous list using a ResNet-50 encoder. In this setting, only previous guidance is used (i.e., $K$=1, without ensemble).
List Length ($N$) Base. 1 2 4 8 12 20
Pascal92 71.9 71.7 71.7 72.4 72.7 72.5 71.9
Pascal183 72.5 72.7 72.9 73.4 73.8 73.5 73.3
4.3.3 Upper Bound Number for Random Selection
To generate reliable and diverse pseudo-labels, we proposed a strategy that randomly selects $k$ models (ranging from 1 to $K$) for each iteration. In this strategy, we explore the performance changes regarding the upper bound number $K$. Table VIII indicates that including the ensembling cases ($K>1$, i.e., $k$=1 or $k>$1 are randomly selected) improves the performance significantly compared to the case of $K=1$ (i.e., without ensemble). In addition, we observe the best results at $K=3$ and a slight performance drop in settings with $K$ greater than 3. Even for large $K$, a varying number ($k$) of models is selected; however, the proportion of large $k$ values increases with $K$. This ensures consistent pseudo-labels but reduces their diversity, potentially degrading performance, as mentioned in Dual Teacher [31]. In conclusion, we select $K=3$ because it adequately satisfies the diversity and reliability requirements of the pseudo-labels.
TABLE VIII: Ablation study on the efficacy of the upper bound number $K$ using a ResNet-50 encoder and $N$=8.
Upper Bound Number ($K$) 1 2 3 4 5
Pascal92 72.7 73.1 73.4 73.4 73.1
Pascal183 73.8 74.9 75.4 75.2 75.1
4.3.4 Criteria for Saving Previous Models
As described in Section 3.2, one alternative for storing previous models involves saving the model at regular intervals, a method used in [62, 63, 21]. Thus, we conduct experiments to validate the effectiveness of this approach. As listed in Table IX, although case (b) exhibits slightly better overall performance than the baseline (a), the difference is marginal. This suggests that storing models at short intervals does not address the coupling problem between the teacher and student networks. In contrast, case (c) shows a significant improvement compared to case (a). Although case (c) functions well, it exhibits limited improvements compared to case (d) which utilizes the proposed save criteria, demonstrating the superiority of the proposed approach. In addition, our approach does not require additional hyperparameter searches to determine appropriate intervals, thereby reducing unnecessary training costs.
TABLE IX: Ablation study regarding the efficacy of the save criteria using a ResNet-50 encoder, $N$=8, and $K$=3.
Save Criteria 92 183 366 732 1464
(a) Baseline 71.9 72.5 76.0 77.4 78.7
(b) Every 1 Epoch 71.8 73.5 76.3 77.5 78.6
(c) Every 3 Epochs 72.2 74.7 76.8 78.0 78.6
(d) On Best Epochs (PrevMatch) 73.4 75.4 77.5 78.6 79.3
Figure 3: Qualitative segmentation results on (a) Pascal VOC with a 92-label partition using a ResNet-50 encoder and (b) Cityscapes with a 1/16 label partition using a ResNet-101 encoder.
4.4 Discussion
TABLE X: Class-wise IoU scores for Pascal VOC with a 92-label partition using a ResNet-50 encoder. $\Delta$ indicates the difference between the PrevMatch and baseline UniMatch performances.
backgr.
airplane
bicycle
bird
boat
bottle
bus
car
cat
chair
cow
table
dog
horse
m.bike
person
plant
sheep
sofa
train
tv
UniMatch 91 84 59 89 71 68 92 80 88 8 88 55 85 85 75 80 54 82 33 80 64
PrevMatch 93 84 61 87 71 68 93 84 89 21 89 57 84 86 77 82 51 82 46 84 58
$\Delta$ 2 0 2 -2 0 0 1 4 1 13 1 2 -1 1 2 2 -3 0 13 4 -6
TABLE XI: Class-wise IoU scores for Cityscapes with a 1/8 label partition using a ResNet-101 encoder.
road
sidewalk
build.
wall
fence
pole
t.light
t.sign
veget.
terrain
sky
person
rider
car
truck
bus
train
m.cycle
bicycle
UniMatch 98 82 92 56 60 62 71 79 92 60 95 82 63 95 83 87 79 68 77
PrevMatch 98 84 92 60 63 63 71 80 92 63 95 82 64 95 83 88 81 69 77
$\Delta$ 0 2 0 4 3 1 0 1 0 3 0 0 1 0 0 1 2 1 0
4.4.1 Analysis of Performance and Training Stability
Class-wise IoU Scores and Qualitative Evaluation. Tables X and XI list the category-wise IoU scores. In particular, Table X on Pascal VOC shows that the proposed method achieves notable performance gains for the chair and sofa classes, which were particularly challenging for the UniMatch baseline. In addition, Table XI on the Cityscapes dataset shows that the proposed method achieves the largest performance improvements for the wall, fence, and terrain classes, which are the lowest performing among the 19 classes in UniMatch. In addition to the quantitative results, the qualitative results shown in Fig. 3 corroborate these findings, revealing consistent improvements in the same categories. Thus, these results suggest that utilizing previous knowledge helps prevent the catastrophic forgetting problem described in Section 3.2, even in semi-supervised semantic segmentation scenarios. We further investigate this problem in the subsequent analysis.
Training Stability of Poorly Behaved Classes. To further explore the catastrophic forgetting problem in poorly behaved classes, we visualize changes in terms of pseudo-label accuracy and validation IoU scores during training, as depicted in Fig. 4. In the first row (chair class), the training curve of the baseline pseudo-label accuracy exhibits significant fluctuations, particularly showing a sudden sharp performance drop at approximately 50 epochs. Although the pseudo-label accuracy recovers slightly thereafter, the validation score does not. In the second row (sofa class), the training curve of the baseline exhibits more severe fluctuations and sharper and more drastic performance drops than in the first row. In contrast, the proposed method shows a smoother training curve without significant fluctuations in either category. This indicates that the proposed training procedure aids in achieving stable optimization for poorly behaved classes that suffer from the forgetting problem.
Figure 4: Training curves for the chair and sofa classes, illustrating variations in pseudo-label pixel accuracy and validation IoU scores. The experiment is conducted on the 92-label partition of Pascal VOC.
Training Stability Across Different Label Partitions. Fig. 5 illustrates the effect of the proposed method on the changes in the validation scores throughout the training process. In fewer label settings (92 and 183), the baseline (blue) exhibits significant fluctuations in terms of performance compared to the proposed method (orange). Moreover, when considering the epoch that achieves the best performance, the baseline method struggles to converge consistently across epochs and tends to become trapped in local minima prematurely. This issue is particularly pronounced in scenarios with fewer labels. In contrast, the proposed method consistently converges across epochs without significant fluctuations. Finally, the consistent outperformance of our method over the baseline across almost all training epochs in label partitions suggests that the positive effects of previous guidance outweigh any negative effects. Note that the positive and negative effects refer to the cases (3) and (2), respectively, described in the last paragraph of Section 3.2.
The results presented in this subsection suggest the following implications: (1) Training with fewer labels leads to instability in pseudo-label predictions, particularly for poorly behaved classes. (2) The proposed training procedure aids in stabilizing pseudo-label predictions while mitigating the forgetting problem in poorly behaved classes. (3) In other words, leveraging previous knowledge helps alleviate the issues of tight coupling and catastrophic forgetting, which can hinder stable learning in semi-supervised segmentation settings.
Figure 5: Training curves for different label partition settings on Pascal VOC. The X- and Y-axes represent epochs and validation mIoU, respectively. The square symbol ($\blacksquare$) denotes the epoch with the best performance.
TABLE XII: Comparison of time spent using existing methods with a ResNet-101 encoder. Training time per epoch and GPU memory usage were measured using the same environment (two A6000 GPUs) and hyperparameters, such as batch size, in a 92-label partition. We used the open-sourced code provided by the authors.
Method #Trainable Networks Training Time (m) GPU Memory Usage (G) Pascal
92 183 366
UniMatch 1 7.8 21.9 75.2 77.2 78.8
Diverse Co-T. 3 22.1 40.5 75.4 76.8 79.6
PS-MT 1 24.9 28.6 65.8 69.6 76.6
ours ($K$=1) 1 8.4 22.7 76.4 77.5 79.3
ours ($K$=3) 1 9.1 22.7 77.0 78.5 79.6
4.4.2 Efficiency Evaluation
To investigate the complexities of the proposed method, we measure the training time per epoch and GPU memory usage. In Table XII, Diverse Co-training, based on the co-training approach, slightly outperforms UniMatch, which operates on a single trainable network, in terms of performance. However, Diverse Co-training requires approximately three times more training time and twice as much memory, resulting in limited scalability. Moreover, PS-MT, a dual Mean Teacher-based method, requires training time comparable to that of Diverse Co-training due to its complex components, despite being a single trainable network. In contrast, the proposed method significantly outperforms existing methods, with a slight cost increase. In particular, PrevMatch ($K$=3) outperforms UniMatch by 1.8% in terms of mIoU (92 label setting) while requiring only a slight increase in complexities in training time and GPU usage, suggesting that PrevMatch can efficiently provide diverse pseudo-labels. Therefore, owing to its computational efficiency and the simplicity of its pipeline reusing the same inputs, the proposed method can be easily integrated into any semi-supervised learning method.
5 Conclusion and Future Work
In this work, we introduced the PrevMatch framework, which leverages temporal knowledge obtained during training to efficiently address the issues of tight coupling and confirmation bias that impede stable semi-supervised learning. The main contributions of PrevMatch include revisiting the use of temporal knowledge and maximizing its effectiveness. Specifically, we directly utilized previous models to provide additional pseudo-label guidance, referred to as previous guidance, to the student network. In addition, we developed a highly randomized ensemble strategy that enhances the reliability and diversity of the previous guidance while minimizing the increase in computational complexity. Experiments were conducted on four benchmark semantic segmentation datasets, revealing that PrevMatch significantly outperforms existing state-of-the-art methods across different evaluation protocols. Furthermore, our findings indicate that leveraging temporal knowledge facilitates stable optimization, particularly for classes that exhibit poor performance and fluctuations. Finally, the computational efficiency and compatibility of the proposed method facilitate its seamless integration into recent semi-supervised semantic segmentation methods.
In future work, we will investigate the capability of PrevMatch to address the domain adaptation problem. This problem poses a more challenging self-training task due to domain discrepancies between the labeled and unlabeled images, a commonly encountered problem in real-world applications. In addition, we observed a phenomenon where significant fluctuations in pseudo-label accuracy for several classes negatively affect generalization ability. For instance, pseudo-label accuracy recovers slightly after a sharp drop in performance; however, the validation score does not. Although the proposed method has shown that it can mitigate these issues, a more in-depth investigation is required for scenarios involving many classes and imbalanced distributions, such as the COCO and ADE20K datasets. Therefore, we intend to explore this phenomenon extensively across different classes, in terms of the relationship between training instability and generalization ability.
Acknowledgments
This research was supported by Brain Korea 21 FOUR. This research was also supported by Korea University Grant (K2403371) and Korea TechnoComplex Foundation Grant (R2112653).

### v2

@@
-Revisiting and Maximizing Temporal Knowledge in Semi-supervised Semantic Segmentation
+PrevMatch: Revisiting and Maximizing Temporal Knowledge in Semi-Supervised Semantic Segmentation
 Abstract
@@
 (2) we design a highly randomized ensemble strategy to maximize the effectiveness of the previous guidance.
-Experimental results on four benchmark semantic segmentation datasets confirm that the proposed method consistently outperforms existing methods across various evaluation protocols.
-In particular, with DeepLabV3+ and ResNet-101 network settings, PrevMatch outperforms the existing state-of-the-art method, Diverse Co-training, by +1.6 mIoU on Pascal VOC with only 92 annotated images, while achieving 2.4 times faster training.
-Furthermore, the results indicate that PrevMatch induces stable optimization, particularly in benefiting classes that exhibit poor performance.
-Code is available at https://github.com/wooseok-shin/PrevMatch.
-Index Terms:
-Consistency regularization, semantic segmentation, semi-supervised learning, temporal knowledge
+PrevMatch, a simple yet effective plug-in method, can be seamlessly integrated into existing semi-supervised learning frameworks with minimal computational overhead.
+Experimental results on three benchmark semantic segmentation datasets show that incorporating PrevMatch into existing methods significantly improves their performance.
+Furthermore, our analysis indicates that PrevMatch facilitates stable optimization during training, resulting in improved generalization performance.
 1 Introduction
-Semantic segmentation [1, 2, 3, 4] is a critical task in various computer vision-related applications, such as autonomous driving [5, 6, 7], medical image analysis [8, 9], robotics [10], among others [11, 12, 13], where the goal is to assign a semantic class label to each pixel in an image.
+Semantic segmentation is a critical task in various computer vision-related applications, such as autonomous driving [wang2020deep, xie2021segformer], medical image analysis [fan2020pranet, shin2022comma], and robotics [milioto2019bonnet], where the goal is to assign a semantic class label to each pixel in an image.
 Despite the recent success of supervised learning-based methods in semantic segmentation, obtaining precise pixel-level annotations for supervised learning is extremely time-consuming and expensive.
@@
 The research has focused on developing methods enabling models to learn effectively from a limited number of labeled images along with a large number of unlabeled images.
-In semi-supervised learning, self-training [14, 15, 16, 17, 18, 19, 20, 21] and consistency regularization [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] have become the predominant approaches for utilizing unlabeled data.
+In semi-supervised learning, self-training [lee2013pseudo, arazo2020pseudo, chen2022debiased, zhou2020time, yang2022st++] and consistency regularization [sajjadi2016regularization, laine2016temporal, miyato2018virtual, xie2020unsupervised, chen2021semi, yang2023revisiting] have become the predominant approaches for utilizing unlabeled data.
 In particular, self-training involves generating pseudo-labels using the predictions of the current model at each iteration for unlabeled samples and leveraging them to train the model in conjunction with labeled data.
@@
 Recent studies have focused on designing frameworks that combine self-training and consistency regularization to exploit the strengths of each method.
-However, self-training-based methods still suffer from the confirmation bias problem [17] even when combined with consistency regularization.
-This problem is attributed to the accumulation of pseudo-label errors produced by the model itself, which exacerbate as self-training progresses.
+However, self-training-based methods still suffer from the confirmation bias problem [arazo2020pseudo], where the model becomes overfitted in the wrong direction, even when combined with consistency regularization.
+This problem is attributed to the accumulation of pseudo-label errors produced by the model itself, which are exacerbated as self-training progresses.
 To mitigate this problem, existing methods distinguish the model prediction processes for supervised outputs and pseudo-labels.
 In other words, the supervised outputs and pseudo-labels are obtained from different predictions, respectively.
-Among them, some studies [34, 35, 32] have adopted the weak-to-strong consistency paradigm from the perspective of input separation, which supervises the prediction from a strongly perturbed input using the pseudo-label generated from its weakly perturbed counterpart (Fig.
+Some studies [sohn2020fixmatch, zou2020pseudoseg, zhong2021pixel, yang2023revisiting] have adopted the weak-to-strong consistency paradigm from the perspective of input separation, which supervises the prediction from a strongly perturbed input using the pseudo-label generated from its weakly perturbed counterpart (Fig.
 1a).
-Its success is based on the idea that more reliable pseudo-labels can be derived from weak perturbations, and strong perturbations aid in mitigating confirmation bias [17, 21], broadening the knowledge and unlabeled data space [36, 32, 29], and shifting the model’s decision boundary to low-density regions [37].
-Because of this advantage, the weak-to-strong consistency paradigm has become a fundamental component of the most recent methods.
-Figure 1: Illustration of the frameworks for (a) FixMatch [34] (or PseudoSeg [35]), (b) Mean Teacher-based structure [26, 38, 39, 40], (c) Co-training [28, 29] (cps: cross pseudo supervision), (d) Dual Mean Teacher [30, 31], and (e) the proposed method.
-In (d), the inputs ($x^{s_{cut}},x^{s_{class}}$) indicate the CutMix [41] and ClassMix [42] augmentations used in Dual Teacher [31].
-Another method to obtain different prediction views involves using network perturbation based on a teacher-student structure.
-In particular, Mean Teacher [26] is a representative approach in semi-supervised segmentation, where a teacher network is derived using an exponential moving average (EMA) of the student model’s weights (Fig.
+Figure 1: Illustration of the frameworks for (a) FixMatch [sohn2020fixmatch], (b) Mean Teacher-based structure [tarvainen2017mean], (c) Co-training [chen2021semi, li2023diverse] (cps: cross pseudo supervision), (d) Dual Mean Teacher [liu2022perturbed, na2023switching], and (e) the proposed method.
+In (d), the inputs ($x^{s_{cut}},x^{s_{class}}$) indicate the CutMix [yun2019cutmix] and ClassMix [olsson2021classmix] augmentations used in Dual Teacher [na2023switching].
+Another method to obtain different prediction views involves using network perturbation based on a teacher-student structure [laine2016temporal, tarvainen2017mean, jin2022semi, wang2022semi, wang2023hunting].
+In particular, Mean Teacher [tarvainen2017mean] is a representative approach in semi-supervised segmentation, where a teacher network is derived using an exponential moving average (EMA) of the student model’s weights (Fig.
 1b).
-Although the Mean Teacher generates somewhat different prediction views between the teacher and student, this method is limited by a coupling problem [27]: as training progresses, the teacher and student become tightly linked, and consequently, the teacher’s predictions become similar to those of the student.
-To mitigate the coupling problem and obtain diverse predictions, certain studies [30, 31] have proposed a dual EMA teacher-based framework where two teachers are alternatively updated at every epoch (Fig.
+Although the Mean Teacher generates somewhat different prediction views between the teacher and student, this method is limited by a coupling problem [ke2019dual]: as training progresses, the teacher and student become tightly linked, and consequently, the teacher’s predictions become similar to those of the student.
+To mitigate the coupling problem and obtain diverse prediction views, certain studies [liu2022perturbed, na2023switching] have proposed a dual EMA teacher-based framework where two teachers are alternatively updated at every epoch (Fig.
 1d).
-Among them, PS-MT [30] produces more reliable pseudo-labels by ensembling the predictions of the two teachers.
+Among them, PS-MT [liu2022perturbed] produces more reliable pseudo-labels by ensembling the predictions of the two teachers.
 By contrast, Na et al.
-[31] reported that ensembling predictions can reduce the diversity of pseudo-labels.
-Thus, they proposed a Dual Teacher framework that alternately activates two teachers in each epoch to generate diversified pseudo-labels.
-Although these studies have demonstrated the benefits of using multiple teachers to mitigate the coupling problem, their ability to provide reliable and diverse pseudo-labels simultaneously remains limited.
+[na2023switching] reported that ensembling predictions can reduce the diversity of pseudo-labels.
+Thus, they proposed the Dual Teacher framework that alternately activates two teachers in each epoch to provide diversified pseudo-labels.
+Although these studies have demonstrated the benefits of using multiple teachers to mitigate the coupling problem, their ability to provide reliable and diverse pseudo-labels over the course of training remains limited, as they typically focus on only one aspect.
+In other words, PS-MT improves the reliability of pseudo-labels but lacks diversity, whereas Dual Teacher suffers from the opposite limitation.
 In addition, they have incorporated additional complex components to ensure diversity between the two teachers.
@@
 Further, distinct types of augmentation provided to each teacher (e.g., CutMix or ClassMix) and layer perturbation are used for Dual Teacher.
-Consequently, this complexity can hinder the scalability and compatibility of these approaches with existing semi-supervised methods.
-Instead of using EMA-based teachers, a co-training paradigm [43, 44, 45, 46, 27, 28, 47, 29] has been widely used to expand prediction views (Fig.
+Consequently, this complexity can hinder the scalability and compatibility of these approaches with existing semi-supervised segmentation methods.
+Instead of using EMA-based teachers, a co-training paradigm [qiao2018deep, ouali2020semi, fan2022ucc, ke2019dual, chen2021semi, wang2023conflict] has been widely used to expand prediction views (Fig.
 1c).
 This paradigm involves simultaneously training multiple networks with different initializations in a mutual teaching manner, where each network supervises the others using pseudo-labels generated from its predictions.
-Building on this concept, subsequent studies [45, 29] have demonstrated that increasing the diversity of pseudo-label views from student networks improves their generalization ability.
-For example, this can involve the use of more co-networks, diverse input domains (e.g., RGB and frequency), or different architectures (e.g., CNN and Transformer).
+Building on this concept, a subsequent study [li2023diverse] has demonstrated that increasing the diversity of pseudo-label views from student networks improves their generalization ability.
 Co-training provides diverse pseudo-label guidance with stability and without concerns regarding the coupling problem.
@@
 Specifically, we save several models at specific epochs during training and utilize their predictions as additional guidance, referred to as previous guidance, which acts as a regularizer in conjunction with standard guidance.
-This strategy addresses the coupling problem and reduces the complexity associated with additional training components.
 Second, we design a highly randomized ensemble strategy to maximize the effectiveness of utilizing the previous guidance.
 This approach involves selecting a random number of models from those previously saved and ensembling their predictions using randomized weights.
-This strategy can efficiently provide diverse and reliable pseudo-labels, while avoiding the significant computational complexity inherent in co-training approaches.
-Extensive experiments conducted across various evaluation protocols on the PASCAL, Cityscapes, COCO, and ADE20K datasets reveal that the proposed PrevMatch significantly outperforms existing methods.
-In particular, compared to Diverse Co-training, the current state-of-the-art method, PrevMatch achieves a +1.6 mIoU improvement on Pascal VOC with 92 labels while accelerating training by 2.4 times.
-In addition to quantitative evaluations, ablation studies and analyses of the proposed components are conducted to explain the success of PrevMatch.
+These strategies can efficiently provide diverse and reliable pseudo-labels while avoiding the complexities inherent in dual EMA- and co-training-based approaches.
+Extensive experiments conducted across various evaluation protocols on the Pascal, Cityscapes, and ADE20K datasets reveal that the proposed PrevMatch significantly improves the performance of different baseline methods.
+In particular, PrevMatch achieves a 0.9-5.4% mean IoU improvement on Pascal VOC with a 92-label setting.
 2 Related Work
-2.1 Semi-supervised Learning
-The central challenge in semi-supervised learning lies in extracting additional supervision for unlabeled samples.
-Previous studies can be categorized into two main approaches: self-training and consistency regularization.
-Self-training-based methods [14, 15, 16, 17, 18, 19, 20] assign pseudo-labels to unlabeled data using the ongoing model predictions and integrate them with labeled data to retrain the model.
-Consistency regularization methods [22, 23, 24, 25, 34, 48, 26, 43, 27, 49] enforce a model to produce invariant predictions for different perturbations, such as input and network perturbations.
-In terms of input perturbations [22, 23, 24, 25, 34, 48], the $\Pi$-model [50, 22, 23] applies simple stochastic perturbations, while VAT [24] applies adversarial perturbations to unlabeled samples.
-Subsequently, UDA [25] expands the perturbation pool using RandAugment [51] to provide diverse augmented samples.
-Furthermore, FixMatch [34], a standard approach in recent semi-supervised learning, adopts a weak-to-strong paradigm where the prediction for a strongly perturbed input is supervised by pseudo-labels generated from the prediction for a weakly perturbed input.
-Network perturbation, another approach to enforcing consistency regularization, has also been widely investigated [23, 26, 43, 27, 49].
-Temporal Ensembling [23] and Mean Teacher [26] methods leverage earlier predictions and weights of the model, respectively, to obtain different prediction views.
-In addition, co-training-based methods [43, 27, 49] adopt multiple networks with different initializations and enforce consistency regularization between their predictions.
-Among the numerous studies discussed, the weak-to-strong paradigm, Mean Teacher structure, and co-training approaches have been widely used to advance semi-supervised semantic segmentation.
 2.1 Semi-supervised Semantic Segmentation
-Owing to the detailed nature of the segmentation task, which involves numerous pixels and multiple classes within an image, semantic segmentation requires more time-consuming and costly efforts to annotate labels and demands sophisticated approaches for accurate performance.
-Thus, recent research has increasingly focused on developing methods for semi-supervised semantic segmentation, based on previous achievements in semi-supervised learning.
-Early studies [52, 53] employed a GAN framework [54] to generate additional supervision for unlabeled images by distinguishing between pseudo-labels and manual labels.
 Recent studies have focused on developing improved methods based on weak-to-strong consistency, Mean Teacher, and co-training paradigms.
-PseudoSeg [35] adopts a weak-to-strong paradigm based on a single network.
-CPS [28, 45] and GCT [55] employ two networks for co-training and demonstrate that co-training outperforms the Mean Teacher approach.
+PseudoSeg [zou2020pseudoseg] adopts a weak-to-strong paradigm based on a single network.
+CPS [chen2021semi] and GCT [ke2020guided] employ two networks for co-training and demonstrate that co-training outperforms the Mean Teacher approach.
+UniMatch [yang2023revisiting] introduces two key enhancements to FixMatch: dual-stream image-level perturbation and feature-level perturbation strategies.
 Subsequent studies have integrated the weak-to-strong paradigm (i.e., input perturbation) into the Mean Teacher or co-training approaches (i.e., network perturbation), demonstrating performance improvements.
-A stream of research that combines the weak-to-strong paradigm with the Mean Teacher structure has proposed techniques such as advanced data augmentation [38, 56, 39, 40], prototype learning [57], curriculum learning [58], symbolic reasoning [59], and dual Mean Teacher [30, 31].
-Among them, PS-MT [30] and Dual Teacher [31] methods implement a dual EMA teacher-based framework to mitigate the coupling problem between the teacher and student models, with two teachers alternately updating each epoch based on the EMA of the student’s weights.
-However, their pipelines inevitably involve complex components to resolve the coupling problem, limiting the scalability and compatibility of these approaches.
-Moreover, regarding pseudo-label generation, PS-MT ensembles the predictions of two teachers to improve reliability, whereas Dual Teacher alternately uses the predictions of each teacher to enhance diversity.
-However, neither method satisfies both reliability and diversity.
-In other words, PS-MT improves the reliability of pseudo-labels but lacks diversity, whereas Dual Teacher suffers from the opposite limitation.
-Another stream of research, combining the weak-to-strong paradigm with co-training, has developed techniques such as a shared backbone with multiple heads [46], conservative-progressive learning [60], and increasing the diversity of co-training [29].
-Among them, Li et al.
-[29] investigated the working mechanism of co-training and discovered that providing distinct pseudo-label views improves generalization ability.
-To this end, they proposed Diverse Co-training, which incorporates variations in the input domains (RGB and frequency) and architectures (CNN and Transformer).
-However, in diverse co-training, the number of whole networks to train increases as pseudo-label views increase.
-In this study, instead of using Mean Teacher or co-training approaches, we revisit the utilization of temporal knowledge and efficiently expand pseudo-label views by maximizing the use of previous models.
-The proposed PrevMatch framework can be seen as simplifying and extending [30, 31] while also enhancing the efficiency of [29].
-Specifically, we eliminate the complex components used in [30, 31], such as distinct augmentation types for each epoch, layer/adversarial feature perturbations, EMA teachers, and a new loss function.
+A stream of research that combines the weak-to-strong paradigm with the Mean Teacher structure has proposed techniques such as advanced data augmentation [hu2021semi, zhao2023augmentation, zhao2023instance], prototype learning [xu2022semi], curriculum learning [ma2023enhanced], symbolic reasoning [liang2023logic], and dual Mean Teacher [liu2022perturbed, na2023switching].
+Another stream of research, combining the weak-to-strong paradigm with co-training, has developed techniques such as a shared backbone with multiple heads [fan2022ucc], conservative-progressive learning [fan2023conservative], and increasing the diversity of co-training [li2023diverse].
+The proposed PrevMatch framework can be seen as simplifying and extending PS-MT [liu2022perturbed] and Dual Teacher [na2023switching] while also enhancing the efficiency of Diverse Co-training [li2023diverse].
+Specifically, we eliminate the complex components used in PS-MT and Dual Teacher, such as distinct augmentation types for each epoch, layer/adversarial feature perturbations, EMA teachers, and a new loss function.
 Instead, PrevMatch reuses the previous models and the weakly perturbed input used in the standard flow, thereby improving the simplicity and compatibility of the overall framework.
 Furthermore, we provide reliable and diverse pseudo-labels to the student network through a highly randomized ensemble strategy.
-Moreover, although both PrevMatch and co-training provide diverse pseudo-label views, PrevMatch operates on a single trainable network with fixed previous models.
+Moreover, although both PrevMatch and Diverse Co-training provide diverse pseudo-label views, PrevMatch operates on a single trainable network with fixed previous models.
 This facilitates greater efficiency in terms of computational and memory costs.
@@
 In the context of semi-supervised learning, several studies have leveraged temporal knowledge obtained from previous training stages.
-Temporal Ensembling [23] accumulates predictions for unlabeled samples across different epochs using the EMA approach and enforces consistency between the current and EMA predictions.
-Mean Teacher [26] averages the model weights across the training steps using the EMA approach, producing a more stable teacher network.
-As a further extension, a dual EMA teacher-based framework was proposed [30, 31], where two teachers are alternately updated at each epoch.
-Moreover, TC-SSL [20] measures the time-consistency (TC) scores of individual samples across training epochs and selects unlabeled samples with higher TC scores for consistency learning.
-This approach assumes that time-consistent predictions are typically accurate.
-Similar to TC-SSL, ST++ [21] proposes a selective re-training scheme that selects more reliable samples based on a stability score.
-The stability score based on the mIoU metric is derived from comparisons between the predictions of the final model and those of several checkpoints obtained during training.
-In addition to semi-supervised learning, Feng et al.
-[61] proposed a temporal consistency framework that adopts two temporary teachers to learn instance temporal consistency in representation learning.
+Temporal Ensembling [laine2016temporal] accumulates predictions for unlabeled samples across different epochs using the EMA approach and enforces consistency between the current and EMA predictions.
+Mean Teacher [tarvainen2017mean] averages the model weights across the training steps using the EMA approach, producing a more stable teacher network.
+PS-MT [liu2022perturbed] and Dual Teacher [na2023switching], inheriting the spirit of the Mean Teacher method, aim to further exploit temporal knowledge by adopting two EMA teachers.
+Moreover, TC-SSL [zhou2020time] and ST++ [yang2022st++] methods implement a filtering mechanism to exclude less-informative unlabeled samples using specific scoring criteria measuring temporal consistency.
+In contrast to existing methods, we directly utilize previous models and generate pseudo-label guidance from their predictions without relying on a complex pipeline of dual EMA-based methods or a filtering mechanism.
 3 Method
@@
 $L=\frac{1}{2}(L_{s}+L_{u}).$ (1)
-Figure 2: Overall framework of PrevMatch (based on UniMatch).
-FP indicates the feature perturbation used in UniMatch.
-$g$ and $h$ are the encoder and decoder of the entire network $f$, respectively.
-When the maximum length of the previous list is exceeded, the oldest model is replaced with the new model using a first-in-first-out (FIFO) approach.
-The red and blue arrows denote standard and proposed guidance flows, respectively.
-  1. 1.
+  1. 1)
     Supervised Flow In supervised flow, a segmentation network $f$ receives a labeled image $x^{l}$ and generates the corresponding predicted class distribution $p^{l}$.
@@
 This can be formulated as follows.
-    $L_{s}=\sum_{i,j}H(y^{l}_{ij},p^{l}_{ij}),$ (2)
-    where $i$ and $j$ indicate pixel indices.
-  2. 2.
-    Unsupervised – Standard Flow Following the success of the weak-to-strong consistency paradigm popularized by FixMatch [34], most semi-supervised segmentation methods have adopted this paradigm.
-Concretely, two perturbed images (i.e., $x^{w}$ and $x^{s}$) are obtained by applying weak and strong augmentations to an unlabeled image $x^{u}$.
+    $L_{s}=\sum H(y^{l},p^{l}).$ (2)
+  2. 2)
+    Unsupervised – Standard Flow Following the success of the weak-to-strong consistency paradigm popularized by FixMatch [sohn2020fixmatch], most semi-supervised segmentation methods have adopted this paradigm.
+Concretely, two perturbed images ($x^{w}$ and $x^{s}$) are obtained by applying weak and strong augmentations to an unlabeled image $x^{u}$.
 The network $f$ receives the two perturbed images and outputs the predicted class distributions $p^{w}$ and $p^{s}$.
@@
 1e (red arrow).
-This consistency term can be formulated as follows.
-    $C(p^{w},p^{s})=\sum_{i,j}\mathbb{1}(max(p^{w}_{ij})\geq\tau)H(y^{w}_{ij},p^{s}_{ij}),$ (3)
+This consistency term can be formulated as:
+    $L_{u(standard)}=\sum\mathbb{1}(max(p^{w})\geq\tau)H(y^{w},p^{s}),$ (3)
     where $\mathbb{1}(\cdot)$ is an indicator function, and $\tau$ is a confidence threshold used to ignore noise in a pseudo-label.
-Furthermore, the UniMatch framework [32] introduced two contributions based on FixMatch: the dual stream image-level perturbation and the feature-level perturbation strategies, to expand the perturbation space, thereby achieving significant improvements.
-Therefore, we select the UniMatch framework as our baseline.
-In particular, two strongly perturbed images ($x^{s_{1}}$ and $x^{s_{2}}$) are derived by randomly applying a strong augmentation pool to the same image, $x^{u}$.
-Subsequently, their corresponding predictions ($p^{s_{1}}$ and $p^{s_{2}}$) are generated through the network $f$.
-Regarding the feature-level perturbation strategy, the prediction is obtained as follows: $p^{fp}=h(\text{Dropout}(g(x^{w})))$, where $g$ and $h$ are the encoder and decoder of the entire network $f$, respectively.
-Ultimately, three predictions are simultaneously supervised by a common pseudo-label derived from a weak view (red arrow in Fig.
-2).
-This can be formulated as follows.
-    $L_{u(standard)}=C(p^{w},p^{s_{1}})+C(p^{w},p^{s_{2}})+C(p^{w},p^{fp}).$ (4)
-  3. 3.
+  3. 3)
     Unsupervised – Proposed Flow To obtain additional pseudo-label guidance, the same weakly perturbed image, $x^{w}$, used in the standard branch, is fed into the previous model branch.
@@
 As illustrated by the blue arrow in Fig.
-2, the previous guidance functions as additional regularizers, supervising the three predictions.
+1e, the previous guidance functions as additional regularizers, supervising the strongly perturbed prediction.
 This can be formulated as follows.
-    $L_{u(prev)}=C(p^{w^{\prime}},p^{s_{1}})+C(p^{w^{\prime}},p^{s_{2}})+C(p^{w^{\prime}},p^{fp}).$ (5)
+    $L_{u(prev)}=\sum\mathbb{1}(max(p^{w^{\prime}})\geq\tau)H(y^{w^{\prime}},p^{s}).$ (4)
 Finally, the total loss of the unsupervised flow is defined by combining the standard and proposed flow losses as follows:
@@
 where $\lambda$ denotes the weight of the proposed flow.
-3.2 Previous Guidance: Revisiting Temporal Knowledge
-To address the coupling problem between student and teacher networks, we revisit the utilization of temporal knowledge that can be obtained during the training process.
-In the literature, Temporal Ensembling [23] and Mean Teacher [26] methods average the temporal knowledge of a model in terms of its predictions and weights, respectively.
-PS-MT [30] and Dual Teacher [31], inheriting the spirit of the Mean Teacher [26] method, aim to further exploit temporal knowledge by adopting two EMA teachers.
-TC-SSL [20] and ST++ [21] methods implement a filtering mechanism to exclude less-informative unlabeled samples using specific scoring criteria measuring temporal consistency.
-In contrast to existing methods, we directly utilize previous models and generate pseudo-label guidance from their predictions without relying on a complex pipeline of dual EMA-based methods or a filtering mechanism.
+3.2 Previous Guidance
+To address the coupling problem between student and teacher networks, we directly utilize previous models and generate pseudo-label guidance from their predictions, instead of using Mean Teacher or co-training approaches.
 In particular, we store multiple models at different epochs that meet the specified criteria during training and use their pseudo-labels as additional guidance, referred to as previous guidance.
@@
 Then, this model processes the same weakly perturbed image, $x^{w}$, used in the standard flow and produces predictions $p^{w^{\prime}}$.
-The previous guidance is obtained by: $y^{w^{\prime}}\text{= argmax}(p^{w^{\prime}})$, and it supervises three predictions according to Eq.
+The previous guidance is obtained by: $y^{w^{\prime}}\text{= argmax}(p^{w^{\prime}})$, and it supervises the strongly perturbed prediction according to Eq.
 (4).
-In this way, we produce diverse pseudo-labels, as in prior studies [30, 31, 28, 29], by leveraging different previous models encompassing varied perspectives of temporal knowledge, without complex additional components or heavy computational burden.
-Save Criteria.
+In this way, we produce diverse pseudo-labels, as in prior studies [liu2022perturbed, na2023switching, chen2021semi, li2023diverse], by leveraging different previous models encompassing varied perspectives of temporal knowledge, without complex additional components or heavy computational burden.
+3.2.1 Save Criteria.
 In this approach, storing the appropriate previous models is crucial for generating diverse and reliable pseudo-labels.
-Some studies [62, 63, 21] that employ intermediate models for ensembling or filtering in image recognition and segmentation save the model at regular intervals (e.g., every 20 epochs in a total of 100 epochs).
-By contrast, we save the model when it achieves the best performance on the validation set to ensure the stability of the previous guidance.
-The weights and performances of the neural network can vary significantly during the optimization process, and excessively large fluctuations may increase negative impacts.
-These negative impacts can be exacerbated by label scarcity and a pronounced class imbalance in semi-supervised semantic segmentation compared to standard image recognition.
-In addition, the periodic saving method requires additional hyperparameter searches to determine appropriate intervals.
-Therefore, we adopt our saving approach to achieve stability and simplicity (refer to Table IX for related experiments).
+Some studies [huang2017snapshot, wang2022efficient] that employ intermediate models for ensembling in image recognition save the model at regular intervals (e.g., every 20 epochs in a total of 100 epochs).
+In contrast to standard image recognition, the weights and performances of the neural network can vary significantly during the optimization process in semi-supervised segmentation (i.e., our case) due to label scarcity and a pronounced class imbalance.
+Therefore, we save the model when it achieves the best performance on the validation set to ensure the stability of the previous guidance.
+In addition, this approach does not require additional hyperparameter searches, such as determining appropriate intervals.
+Therefore, we adopt our saving approach to achieve stability and simplicity.
 The efficacy of the previous guidance can be intuitively explained as follows.
@@
 In contrast, given the standard and previous guidance, four scenarios can be considered based on whether they are correct or incorrect (standard-previous): (1) correct-correct, (2) correct-incorrect, (3) incorrect-correct, and (4) incorrect-incorrect.
-Through case (3), the network receives an additional opportunity to be guided in the right direction, away from the wrong one.
-Although model training may be hindered through case (2), leading to significant fluctuations, we empirically demonstrate that the positive effects of the proposed method outweigh the negative effects (which is further supported by the analysis of training stability in Section 4.4.1).
-In addition, previous guidance can help mitigate the network’s catastrophic forgetting problem [64, 65], where previously learned knowledge is forgotten when acquiring new knowledge.
-In particular, this phenomenon can be more pronounced in self-training and class-imbalance scenarios (i.e., in our scenario) due to the lack of labeled data, potentially leading to significant performance fluctuations in poorly behaved classes [19].
-In semi-supervised [20] and representation learning [61], some studies have demonstrated that utilizing temporal knowledge helps mitigate the catastrophic forgetting problem and stabilizes training.
-Based on this fact, we explore the learning stability of the proposed method in Section 4.4.1.
+Through case (3), the network receives an additional opportunity to be guided in the right direction, away from the wrong one (Refer to Sec.
+8.2 in the supplementary material for further discussion).
 3.3 Maximizing Efficacy of Previous Guidance
@@
 These techniques have been widely used in various domains as a promising method for improving performance.
-In particular, several studies [62, 63] in the field of image recognition have demonstrated that ensembles of multiple intermediate models obtained during training also improve prediction accuracy and diversity.
+In particular, several studies [huang2017snapshot, wang2022efficient] in the field of image recognition have demonstrated that ensembles of multiple intermediate models obtained during training also improve prediction accuracy and diversity.
 Therefore, we utilize network ensembling to improve the reliability of the previous guidance.
@@
 To this end, given the list that includes $N$ previous models, we randomly select $K$ ($K\leq N$) models for each iteration to mitigate the increase in computational complexity while ensuring pseudo-label diversity.
-However, this approach, which always ensembles $K$ models, may not guarantee the diversity of pseudo-labels, as noted in Dual Teacher [31].
+However, this approach, which always ensembles $K$ models, may not guarantee the diversity of pseudo-labels, as noted in Dual Teacher [na2023switching].
 This problem can worsen for large values of $K$.
@@
   • Random Selection: For each iteration, we randomly select a varying number of teachers, $k$, ranging from 1 to $K$.
-For example, $k$=1,2, or 3 can be selected for each iteration when $K$=3.
+For example, kk=1,2, or 3 can be selected for each iteration when KK=3.
 With this approach, selecting a large $k$ tends to yield consistent pseudo-labels, enhancing reliability, while a small $k$ contributes to increased diversity in the pseudo-labels.
@@
 This approach explores all combinations of previous guidance in a continuous space, expanding the original pseudo-label space beyond a simple average.
-In this way, the proposed randomized ensemble strategy can improve the reliability and diversity of the previous guidance while mitigating the increase in computational complexity.
+In this way, the proposed randomized ensemble strategy probabilistically generates pseudo-labels with varying degrees of reliability and diversity across the training process, thereby enabling broader and more flexible exploration of the pseudo-label space.
 4 Experiments
-4.1 Experimental Configuration
-4.1.1 Datasets
-PASCAL VOC [66] is a widely used benchmark dataset in semantic segmentation, comprising 21 distinct categories (20 object types and a background class).
-This dataset contains 10,582 images designated for training purposes, segmented into two subsets based on annotation quality: 1,464 images constitute the high-quality subset, characterized by detailed annotations, whereas the remaining 9,118 images form the coarse subset, featuring less detailed annotations.
-This study follows established protocols for a fair comparison [28, 67, 32, 39].
-In particular, three training protocols are considered based on the criteria for selecting labeled images.
-  • Original: A protocol whereby labeled images are exclusively sourced from the high-quality subset.
-  • Blended: A protocol that entails a random selection of labeled images from the total dataset.
-  • Priority: A protocol where the selection of labeled images is first derived from the high-quality subset; if not sufficient, it is complemented by additional images from the coarse subset.
-Cityscapes dataset [68], tailored for the semantic analysis of urban street scenes, comprises 2,975 high-resolution images for training and 500 images for validation, primarily focusing on 19 categories within urban environments.
-Consistent with previous studies [21, 32, 29], we evaluate this dataset using various label partitions, specifically 1/16, 1/8, 1/4, and 1/2 of the total number of labels.
-COCO dataset [69], notable for its complexity and scale, comprises 118k training and 5k validation images.
-This dataset features dense annotations across 81 classes set in various indoor and outdoor scenarios.
-Considering the performance plateaus observed in datasets such as Pascal and Cityscapes and the higher number of classes, the COCO dataset has emerged as a practical and valuable benchmark for evaluating advanced algorithms in semi-supervised segmentation.
-Following existing studies [32], we validate the proposed method using 1/256, 1/128, and 1/64 label partitions.
-ADE20K dataset [70], containing more diverse scenes and a greater number of classes (150 categories) than COCO, comprises 20,210 images for training and 2,000 images for validation.
-Using the label partitions from existing studies [33], we evaluate the proposed method on 1/128, 1/64, and 1/32 partitions.
+4.1 Datasets
+(1) Pascal VOC dataset [everingham2010pascal] contains 10,582 training images, segmented into two subsets based on annotation quality: 1,464 images constitute the high-quality subset, whereas the remaining 9,118 images form the coarse subset.
+In particular, two training protocols are considered based on the criteria for selecting labeled images: Original protocol whereby labeled images are exclusively sourced from the high-quality subset.
+Blended protocol that entails a random selection of labeled images from the total dataset.
+(2) Cityscapes dataset [cordts2016cityscapes], tailored for the semantic analysis of urban street scenes, comprises 2,975 high-resolution images for training and 500 images for validation, primarily focusing on 19 categories within urban environments.
+(3) ADE20K dataset [zhou2017scene], containing more diverse scenes and a greater number of classes (150 categories) than the COCO dataset [lin2014microsoft], comprises 20,210 images for training and 2,000 images for validation.
 In all protocols, training images not selected as labeled images are utilized as unlabeled images.
-4.1.2 Architecture and Implementation Details
-Consistent with existing literature, we employ ResNet-50 and ResNet-101 backbones [71] as encoders for the Pascal VOC, Cityscapes, and ADE20K datasets.
-For the COCO dataset, Xception-65 [72] is adopted as the backbone.
-We use DeepLabV3+ [3] as the segmentation head and set its output stride to 16 for efficient training.
-For the training setups, each mini-batch comprises 8 labeled and 8 unlabeled images.
-The proposed method is trained for 80, 240, 30, and 40 epochs for the PASCAL, Cityscapes, COCO, and ADE20K datasets, respectively, using the SGD optimizer.
-The learning rates are initially set to 0.001, 0.005, 0.004, and 0.004 for these datasets, respectively, and are managed using a polynomial learning rate scheduler.
-Moreover, we use 321/513, 801, 513, and 513 random crops for these datasets, respectively.
-For image augmentation, we use common weak (e.g., resize, crop, and flip) and strong (e.g., color transformations, grayscale, cutmix, and blur) data augmentations, as in UniMatch [32].
+4.2 Implementation Details
+We select various methods that potentially exhibit confirmation bias and coupling problems as baselines, which use the input perturbation method based on a single network.
+For the four baseline methods (FixMatch, UniMatch, AugSeg, and MPMC), we employ ResNet-50/101 [he2016deep] as the encoder and DeepLabV3+ [chen2018encoder] as the segmentation head.
+For UniMatchV2, we use the DINOv2 [oquab2023dinov2] encoder and the DPT [ranftl2021vision] segmentation model.
+Specifically, UniMatch and UniMatchV2 are the primary baselines in our study.
+For these two baselines, each mini-batch comprises 8 labeled and 8 unlabeled images.
+Furthermore, they are trained for 80, 240/180, and 40/60 epochs for the PASCAL, Cityscapes, and ADE20K datasets, respectively.
+The learning rates are initially set to 1e-3/5e-6, 5e-3/5e-6, and 4e-3/5e-6 for these datasets, respectively, and are managed using a polynomial learning rate scheduler.
+Moreover, we use 321/518, 801/798, and 513/518 random crops for these datasets, respectively.
+For image augmentation, we use common weak (e.g., resize, crop, and flip) and strong (e.g., color transformations, grayscale, cutmix, and blur) data augmentations, as in UniMatch [yang2023revisiting].
 For the hyperparameters of standard consistency flow, $\tau$ is set to 0 for Cityscapes and 0.95 for the other datasets.
-The dropout rate for feature perturbation is set to 0.5.
-For the hyperparameters of PrevMatch, $\tau$ is set to 0.9 for Pascal, 0 for Cityscapes, and 0.95 for COCO and ADE20k.
-The maximum length ($N$) of the previous list is set to eight for Pascal and Cityscapes and five for COCO and ADE20K.
+For the hyperparameters of PrevMatch, $\tau$ is set to 0.9 for Pascal, 0 for Cityscapes, and 0.95 for ADE20k.
+The maximum length ($N$) of the previous list is set to eight for Pascal and Cityscapes and six for ADE20K.
 The upper bound number $K$ for the random selection is set to three.
-For weight $\lambda$, as previous guidance should correct the model before it becomes overfitted in the wrong direction (i.e., confirmation bias [17]) during the middle of training, setting an appropriate weight $\lambda$ is crucial.
-Additionally, as the initial model typically exhibits poor performance, we implement a warmup schedule for the weight of the proposed flow, similar to the commonly used polynomial learning rate scheduler that includes a warmup phase.
-Experiments are conducted in the following environments: UBUNTU 20.04, Python 3.10.4, PyTorch 1.12.1, CUDA 11.3, and NVIDIA RTX 3090Ti or A6000 GPUs.
-The training is performed using one GPU for Pascal and two GPUs for the other datasets.
-TABLE I: Comparison with state-of-the-art methods on the Original protocol of Pascal VOC dataset.
-The numeric values in the header (e.g., 92) represent the number of labeled images used for training.
-All methods are trained using ResNet-50/101 and DeepLabV3+.
-The number of trainable whole networks and input resolution are reported.
-The values of PrevMatch are averaged over three runs.
-The evaluation metric is the mean IoU (%).
-Pascal [Original set] Encoder #Trainable Networks Resolution # Labeled images (Total: 10582)
-92 183 366 732 1464
-Supervised Baseline R-50 $\times$1 5132 44.0 52.3 61.7 66.7 72.9
-PseudoSeg [35] [ICLR’21] R-50 $\times$1 5132 54.9 61.9 64.9 70.4 71.0
-PC2Seg [73] [ICCV’21] R-50 $\times$1 5132 56.9 64.6 67.6 70.9 72.3
-CPCL [60] [TIP’23] R-50 $\times$2 5122 61.9 67.0 72.1 74.3 -
-AugSeg [39] [CVPR’23] R-50 $\times$1 5122 64.2 72.2 76.2 77.4 78.8
-UniMatch [32] [CVPR’23] R-50 $\times$1 3212 71.9 72.5 76.0 77.4 78.7
-Dual Teacher [31] [NeurIPS’23] R-50 $\times$1 3212 70.8 74.5 76.4 77.7 78.2
-PrevMatch (ours) R-50 $\times$1 3212 73.4 75.4 77.5 78.6 79.3
-Supervised Baseline R-101 $\times$1 5132 45.1 55.3 64.8 69.7 73.5
-CPS [28] [CVPR’21] R-101 $\times$2 5122 64.1 67.4 71.7 75.9 -
-ReCo [74] [ICLR’22] R-101 $\times$1 3212 64.8 72.0 73.1 74.7 -
-PS-MT [30] [CVPR’22] R-101 $\times$1 5122 65.8 69.6 76.6 78.4 80.0
-ST++ [21] [CVPR’22] R-101 $\times$1 3212 65.2 71.0 74.6 77.3 79.1
-U2PL [67] [CVPR’22] R-101 $\times$1 5122 68.0 69.2 73.7 76.2 79.5
-GTA-Seg [75] [NeurIPS’22] R-101 $\times$2 5132 70.0 73.2 75.6 78.4 80.5
-PCR [57] [NeurIPS’22] R-101 $\times$1 5132 70.1 74.7 77.2 78.5 80.7
-DGCL [76] [CVPR’23] R-101 $\times$1 5132 70.5 77.1 78.7 79.2 81.6
-CCVC [47] [CVPR’23] R-101 $\times$2 5122 70.2 74.4 77.4 79.1 80.5
-iMAS [40] [CVPR’23] R-101 $\times$1 5132 68.8 74.4 78.5 79.5 81.2
-AugSeg [39] [CVPR’23] R-101 $\times$1 5122 71.1 75.5 78.8 80.3 81.4
-UniMatch [32] [CVPR’23] R-101 $\times$1 3212 75.2 77.2 78.8 79.9 81.2
-ESL [58] [ICCV’23] R-101 $\times$1 5132 71.0 74.1 78.1 79.5 81.8
-LogicDiag [59] [ICCV’23] R-101 $\times$1 5132 73.3 76.7 77.9 79.4 -
-LogicDiag + MKD [59] [ICCV’23] R-101 $\times$2 5132 74.7 77.2 78.4 80.1 -
-Diverse Co-T. (2-cps) [29] [ICCV’23] R-101 $\times$2 3212 74.8 77.6 79.5 80.3 81.7
-Diverse Co-T. (3-cps) [29] [ICCV’23] R-101 $\times$3 3212 75.4 76.8 79.6 80.4 81.6
-PrevMatch (ours) R-101 $\times$1 3212 77.0 78.5 79.6 80.4 81.6
-TABLE II: Comparison with state-of-the-art methods on the Blended protocol of Pascal VOC dataset.
-All methods are trained using ResNet-50 and DeepLabV3+.
-The fractional values indicate the ratio of labeled images used for training.
-Pascal [Blended set] Resolution 1/16 1/8 1/4
-Supervised Baseline 5132 62.4 68.2 72.3
-Mean Teacher [26] [NeurIPS’17] 5122 66.8 70.8 73.2
-CCT [44] [CVPR’20] 5122 65.2 70.9 73.4
-GCT [55] [ECCV’20] 5122 64.1 70.5 73.5
-CutMix-Seg [38] [BMVC’20] 5122 68.9 70.7 72.5
-CAC [77] [CVPR’21] 3202 70.1 72.4 74.0
-CPS [28] [CVPR’21] 5122 72.0 73.7 74.9
-UCC [46] [CVPR’22] 5122 74.1 74.8 76.4
-PS-MT [30] [CVPR’22] 5122 72.8 75.7 76.4
-UniMatch [32] [CVPR’23] 3212 74.5 75.8 76.1
-AugSeg [39] [CVPR’23] 5122 74.7 76.0 77.2
-iMAS [40] [CVPR’23] 5132 74.8 76.5 77.0
-CCVC [47] [CVPR’23] 5122 74.5 76.1 76.4
-PrevMatch (ours) 3212 75.6 76.4 76.3
-5132 76.0 77.1 77.6
-TABLE III: Comparison with state-of-the-art methods on the Priority protocol of Pascal VOC dataset. All methods are trained using ResNet-101 and DeepLabV3+.
-Pascal [Priority set] Resolution 1/16 1/8 1/4
-Supervised Baseline 5132 70.6 75.0 76.5
-U2PL [67] [CVPR’22] 5122 77.2 79.0 79.3
-UniMatch [32] [CVPR’23] 5132 80.9 81.9 80.4
-AugSeg [39] [CVPR’23] 5122 79.3 81.5 80.5
-Dual Teacher [31] [NeurIPS’23] 5132 80.1 81.5 80.5
-PrevMatch (ours) 5132 81.4 81.9 80.8
+For weight $\lambda$, as previous guidance should correct the model before it becomes overfitted in the wrong direction (i.e., confirmation bias [arazo2020pseudo]) during the middle of training, setting an appropriate weight $\lambda$ is crucial.
+Additionally, as the initial model typically exhibits poor performance, we implement a warmup schedule for the weight of the proposed flow, similar to the commonly used learning rate decay scheduling that includes a warmup phase.
+Refer to Sec.
+7.6 in the supplementary material for further details.
+Pascal [Original set] Encoder 92 183 366 732 1464
+AugSeg [zhao2023augmentation] [CVPR’23] RN-50 64.2 72.2 76.2 77.4 78.8
+Dual Teacher [na2023switching] [NeurIPS’23] RN-50 70.8 74.5 76.4 77.7 78.2
+FixMatch [sohn2020fixmatch] [NeurIPS’20] RN-50 63.8 70.3 73.2 76.8 77.8
++ PrevMatch RN-50 66.1 (+2.3) 73.1 (+2.8) 75.7 (+2.5) 77.8 (+1.0) 78.5 (+0.7)
+UniMatch [yang2023revisiting] [CVPR’23] RN-50 71.9 72.5 76.0 77.4 78.7
++ PrevMatch RN-50 73.4 (+1.5) 75.4 (+2.9) 77.5 (+1.5) 78.6 (+1.2) 79.3 (+0.6)
+U2PL [wang2022semi] [CVPR’22] RN-101 68.0 69.2 73.7 76.2 79.5
+PS-MT [liu2022perturbed] [CVPR’22] RN-101 65.8 69.6 76.6 78.4 80.0
+Diverse Co-T.(3cps) [li2023diverse] [ICCV’23] RN-101 75.4 76.8 79.6 80.4 81.6
+CorrMatch [sun2024corrmatch] [CVPR’24] RN-101 76.4 78.5 79.4 80.6 81.8
+U2PL+ [wang2024using] [IJCV’24] RN-101 69.3 73.4 75.0 77.1 79.5
+AugSeg [zhao2023augmentation] [CVPR’23] RN-101 71.1 75.5 78.8 80.3 81.4
++ PrevMatch RN-101 73.4 (+2.3) 77.3 (+1.8) 80.0 (+1.2) 80.9 (+0.6) 82.0 (+0.6)
+UniMatch [yang2023revisiting] [CVPR’23] RN-101 75.2 77.2 78.8 79.9 81.2
++ PrevMatch RN-101 77.0 (+1.8) 78.5 (+1.3) 79.6 (+0.8) 80.4 (+0.5) 81.6 (+0.4)
+AllSpark [wang2024allspark] [CVPR’24] MiT-B5 76.1 78.4 79.8 80.8 82.1
+SemiVL [hoyer2023semivl] [ECCV’24] CLIP-B 84.0 85.6 86.0 86.7 87.3
+UniMatchV2-S [yang2025unimatch] [TPAMI’25] DINOv2-S 79.0 85.5 85.9 86.7 87.8
++ PrevMatch DINOv2-S 84.4 (+5.4) 86.2 (+0.7) 86.7 (+0.8) 87.5 (+0.8) 88.2 (+0.4)
+UniMatchV2-B [yang2025unimatch] [TPAMI’25] DINOv2-B 86.3 87.9 88.9 90.0 90.8
++ PrevMatch DINOv2-B 87.2 (+0.9) 88.6 (+0.7) 90.5 (+1.6) 90.6 (+0.6) 91.3 (+0.5)
+Table 1: Comparison with state-of-the-art methods on the Original protocol of Pascal VOC. The evaluation metric is the mean IoU (%).
+Pascal [Blended set] Encoder 1/16 1/8 1/4
+Mean Teacher [tarvainen2017mean] RN-50 66.8 70.8 73.2
+PS-MT [liu2022perturbed] RN-50 72.8 75.7 76.4
+CorrMatch [sun2024corrmatch] RN-101 77.6 77.8 78.3
+AllSpark [wang2024allspark] MiT-B5 78.3 80.0 80.4
+FixMatch [sohn2020fixmatch] RN-50 72.5 73.0 74.2
++ PrevMatch RN-50 73.9 74.3 75.1
+Gain ($\Delta$) (+1.4) (+1.3) (+0.9)
+UniMatch [yang2023revisiting] RN-101 76.5 77.0 77.2
++ PrevMatch RN-101 77.8 78.0 78.0
+Gain ($\Delta$) (+1.3) (+1.0) (+0.8)
+UniMatchV2-S [yang2025unimatch] DINOv2-S 82.2 83.6 83.8
++ PrevMatch DINOv2-S 83.3 84.5 84.5
+Gain ($\Delta$) (+1.1) (+0.9) (+0.7)
+Table 2: Comparison with state-of-the-art methods on the Blended protocol of Pascal VOC dataset.
 4.3 Comparison with State-of-the-Art Methods
 PASCAL VOC.
-Three main experiments are conducted on the Original, Blended, and Priority protocols, which are divided according to the criteria for selecting labeled images.
-In Table I, the proposed method outperforms state-of-the-art methods in almost all partitions, even with a single trainable network and a lower resolution.
-In particular, PrevMatch significantly improves performance in settings with fewer labels (92 and 183).
-Compared to Diverse Co-T, PrevMatch exhibits comparable performance in settings with more than 366 labels, but it operates on a single trainable network, indicating that the proposed method is efficient with respect to training costs.
-For Blended and Priority protocols reported in Tables II and III, the proposed method consistently outperforms existing methods.
+Two main experiments are conducted on the Original and Blended protocols.
+In Tab.
+1, applying PrevMatch to each baseline consistently improves performance across all baseline methods.
+Notably, we observe that PrevMatch yields the most improvements in settings with fewer labels (92 and 183).
+In the 92-label setting (ResNet-101), it improves AugSeg, UniMatch, and MPMC by 2.3%, 1.8%, and 1.2%, respectively.
+Furthermore, even on UniMatchV2, which already exhibits strong performance with its transformer structure, PrevMatch further enhances performance.
+Moreover, Tab.
+2 reports results on the Blended protocol.
+PrevMatch consistently improves the performance of all baseline methods.
 Cityscapes.
-Experimental results for the Cityscapes validation set are listed in Table IV.
-Under four label partitions using ResNet-50 and ResNet-101 backbones, the proposed method outperforms previous methods in six out of eight cases.
-Similar to the results reported in Table 1, PrevMatch consistently improves performance in setups with fewer labels.
-However, certain methods (e.g., AugSeg, AEL, and Dual Teacher) perform better than PrevMatch at the 1/2 label setups.
-This can be attributed to the advanced data augmentation techniques used in these methods.
-In other words, these methods could achieve better performance because the 1/2 label partition increasingly resembles a supervised learning setting, where augmentation techniques play a crucial role.
-Thus, we intend to explore the efficacy of the advanced augmentations on the proposed method in future work.
-TABLE IV: Comparison with state-of-the-art methods on the Cityscapes dataset. All methods are trained using ResNet-50/101 and DeepLabV3+.
-Cityscapes Encoder #Trainable 1/16 1/8 1/4 1/2
-Networks (186) (372) (744) (1488)
-Supervised Baseline R-50 $\times$1 63.3 70.2 73.1 76.6
-PS-MT [30] [CVPR’22] R-50 $\times$1 - 75.8 76.9 77.6
-U2PL [67] [CVPR’22] R-50 $\times$1 70.6 73.0 76.3 77.2
-UniMatch [32] [CVPR’23] R-50 $\times$1 75.0 76.8 77.5 78.6
-iMAS [40] [CVPR’23] R-50 $\times$1 74.3 77.4 78.1 79.3
-AugSeg [39] [CVPR’23] R-50 $\times$1 73.7 76.5 78.8 79.3
-CCVC [47] [CVPR’23] R-50 $\times$2 74.9 76.4 77.3 -
-FPL (w/ CPS) [78] [CVPR’23] R-50 $\times$2 74.8 77.3 78.5 -
-Diverse Co-T. (3-cps) [29] [ICCV’23] R-50 $\times$3 - 76.5 77.9 -
-PrevMatch (ours) R-50 $\times$1 75.8 77.8 78.8 79.2
-Supervised Baseline R-101 $\times$1 66.3 72.8 75.0 78.0
-CPS [28] [CVPR’21] R-101 $\times$2 69.8 74.3 74.6 76.8
-AEL [56] [NeurIPS’21] R-101 $\times$1 75.8 77.9 79.0 80.3
-PS-MT [30] [CVPR’22] R-101 $\times$1 - 76.9 77.6 79.1
-U2PL [67] [CVPR’22] R-101 $\times$1 74.9 76.5 78.5 79.1
-PCR [57] [NeurIPS’22] R-101 $\times$1 73.4 76.3 78.4 79.1
-CISC-R [79] [TPAMI’23] R-101 $\times$1 - 75.9 77.7 -
-FPL (w/ AEL) [78] [CVPR’23] R-101 $\times$1 76.6 78.2 78.5 -
-AugSeg [39] [CVPR’23] R-101 $\times$1 75.2 77.8 79.6 80.4
-UniMatch [32] [CVPR’23] R-101 $\times$1 76.6 77.9 79.2 79.5
-ESL [58] [ICCV’23] R-101 $\times$1 75.1 77.2 78.9 80.5
-UPC (w/ U2PL) [80] [ICCV’23] R-101 $\times$1 75.3 77.4 79.0 79.6
-Diverse Co-T. (2-cps) [29] [ICCV’23] R-101 $\times$2 75.0 77.3 78.7 -
-Diverse Co-T. (3-cps) [29] [ICCV’23] R-101 $\times$3 75.7 77.4 78.5 -
-Dual Teacher [31] [NeurIPS’23] R-101 $\times$1 76.8 78.4 79.5 80.5
-PrevMatch (ours) R-101 $\times$1 77.7 78.9 80.1 80.1
-COCO & ADE20K.
-Table V lists results for the large-scale datasets COCO and ADE20K.
-The proposed method consistently improves the performance across all partitions compared to the baseline, UniMatch, suggesting that PrevMatch is also effective on large-scale datasets.
-TABLE V: Evaluation results on the large-scale datasets using Xception-65 for COCO and ResNet-50 for ADE20K.
-Method COCO ADE20K
-1/256 1/128 1/64 1/128 1/64 1/32
-Supervised 28.0 33.6 37.8 7.2 9.9 13.7
-UniMatch [32] 38.9 44.4 48.2 13.6 18.3 23.9
-PrevMatch (ours) 40.2 45.7 48.4 15.4 19.6 24.9
-4.3 Ablation Studies
-In the ablation studies and discussion, the baseline refers to the UniMatch method [32].
-4.3.1 Individual Efficacy of the Proposed Components
-The effects of the individual components are investigated, and the results are presented in Table VI.
-The first row indicates the baseline.
+Experimental results for the Cityscapes validation set are presented in Tab.
+3.
+The results indicate that applying PrevMatch to the four baseline approaches consistently improves their performance across all label partitions.
+Similar to the results reported in Tab.
+1, PrevMatch significantly improves performance in setups with fewer labels.
+In particular, in the 1/16 label setup, PrevMatch improves AugSeg, UniMatch, UniMatchV2-S, and UniMatchV2-B by 1.4%, 1.1%, 1.1%, and 0.8%, respectively.
+Cityscapes 1/16 1/8 1/4 1/2
+(186) (372) (744) (1488)
+ResNet-101
+PS-MT [liu2022perturbed] - 76.9 77.6 79.1
+Diverse Co-T.(3cps) [li2023diverse] 75.7 77.4 78.5 -
+Dual Teacher [na2023switching] 76.8 78.4 79.5 80.5
+CorrMatch [sun2024corrmatch] 77.3 78.5 79.4 80.4
+AugSeg [zhao2023augmentation] 75.2 77.8 79.6 80.4
++ PrevMatch 76.6 78.9 80.3 80.8
+Gain ($\Delta$) (+1.4) (+1.1) (+0.7) (+0.4)
+UniMatch [yang2023revisiting] 76.6 77.9 79.2 79.5
++ PrevMatch 77.7 78.9 80.1 80.1
+Gain ($\Delta$) (+1.1) (+1.0) (+0.9) (+0.6)
+Tansformer-based
+AllSpark [wang2024allspark] 78.3 79.2 80.6 81.4
+SemiVL [hoyer2023semivl] 77.9 79.4 80.3 80.6
+UniMatchV2-S [yang2025unimatch] 80.6 81.9 82.4 82.6
++ PrevMatch 81.7 82.7 83.2 83.2
+Gain ($\Delta$) (+1.1) (+0.8) (+0.8) (+0.6)
+UniMatchV2-B [yang2025unimatch] 82.6 83.3 83.5 83.9
++ PrevMatch 83.4 84.1 84.5 84.8
+Gain ($\Delta$) (+0.8) (+0.8) (+1.0) (+0.9)
+Table 3: Comparison with state-of-the-art methods on the Cityscapes dataset.
+$\dagger$: The results of UniMatchV2 (DINOv2-B) were reproduced with a training resolution of 7562 due to limitations in our available memory.
+ADE20K.
+Tab.
+4 lists results for the large-scale and challenging dataset ADE20K.
+PrevMatch consistently improves baseline performance across all partitions, suggesting that PrevMatch is also effective on the large-scale dataset.
+ADE20K 1/128 1/64 1/32 1/16 1/8
+UniMatch-RN50 [yang2023revisiting] 13.6 18.3 23.9 27.2 30.9
++ PrevMatch 15.4 19.6 24.9 28.3 31.6
+Gain ($\Delta$) (+1.8) (+1.3) (+1.0) (+1.1) (+0.7)
+UniMatchV2-S [yang2025unimatch] 20.9 29.1 34.5 37.0 39.4
++ PrevMatch 22.2 29.8 35.2 37.8 40.0
+Gain ($\Delta$) (+1.3) (+0.7) (+0.7) (+0.8) (+0.6)
+UniMatchV2-B [yang2025unimatch] 25.0 33.7 40.5 42.4 45.1
++ PrevMatch 25.7 35.1 41.4 43.4 46.0
+Gain ($\Delta$) (+0.7) (+1.4) (+0.9) (+1.0) (+0.9)
+Table 4: Evaluation results on the large-scale dataset ADE20K, as reproduced in our experimental setup.
+Previous Simple Random Random w/ UniM. w/ UniM.V2
+Guidance Ensemble Selection Weights 92 92
+- - - - 71.9 79.0
+- - - 72.7 82.2
+- - 72.7 82.7
+- - 73.2 83.8
+- 73.4 84.4
+Table 5: Ablation study of PrevMatch’s components based on UniMatch (ResNet-50) and UniMatch V2 (DINOv2-S) on Pascal VOC with 92 labels. For the ensemble, we set KK=3.
+5 Discussion
+5.1 Ablation Study
+The effects of the individual components are investigated, and the results are presented in Tab.
+5.
+The first row indicates the baselines.
 All components of the proposed method consistently achieve performance gains.
-In particular, previous guidance, which randomly selects one previous model from the previous list for each iteration to generate additional guidance, surpasses the baseline by 0.8% and 1.3% in the 92- and 183-label settings, respectively.
+In particular, previous guidance, which randomly selects one previous model from the previous list for each iteration to generate additional guidance, surpasses the baselines by 0.8% and 3.2% in the 92-label settings, respectively.
 Regarding the number of models ($k$) for the network ensemble, we experiment with simple ensemble (fixed $k$) or random selection (random $k$) strategies.
@@
 In contrast, the results using random $k$ (fourth row) indicate significantly improved performance.
-This implies that a fixed $K$ ensemble improves the reliability of pseudo-labels but limits the diversity, as noted in Dual Teacher [31], whereas a random $K$ strategy can provide reliable and diverse guidance to the model.
+This implies that a fixed $k$ ensemble improves the reliability of pseudo-labels but limits the diversity, as noted in Dual Teacher [na2023switching], whereas a random $k$ strategy can provide reliable and diverse guidance to the model.
 In addition, using random weights for network ensembling contributes to performance gains, indicating that it expands the original pseudo-label space by generating diversified guidance.
-TABLE VI: Ablation study of the components of the proposed method using a ResNet-50 encoder.
-For the ensemble, we set $K$=3.
-The means and standard deviation are also reported based on three runs to validate statistical significance.
-Previous Simple Random Random PASCAL
-Guidance Ensemble Selection Weights 92 183
-- - - - 71.9±0.5 72.5±0.7
-- - - 72.7±0.5 73.8±0.6
-- - 72.7±0.4 74.1±0.5
-- - 73.2±0.4 74.9±0.6
-- 73.4±0.4 75.4±0.5
-4.3.2 Previous List Length
+Refer to Tabs.
+10, 11, 13 and 12 in the supplementary material for additional ablation studies on different protocols and the hyperparameters of each component.
+chair
+sofa
+car
+train
+m.bike
+bicycle
+d.table
+backgr.
+UniMatch 8 33 80 80 75 59 55 91
++ PrevMatch 21 46 84 84 77 61 57 93
+Gain ($\Delta$) $cellcolorgreen!3013 \cellcolorgreen$$\cellcolorgreen!154$ $cellcolorgreen!102 \cellcolorgreen$$\cellcolorgreen!102$
+Table 6: Class-wise IoU scores (Top 8 classes based on gain) for Pascal VOC with a 92-label partition using ResNet-50.
+5.2 Analysis of Performance and Training Stability
+Tab.
+6 lists the category-wise IoU scores and shows that the proposed method achieves notable performance gains for the chair and sofa classes, which were particularly challenging for the UniMatch baseline.
+To further explore these poorly behaved classes, we visualize changes in terms of pseudo-label accuracy and validation IoU scores during training, as depicted in Fig.
+2.
+In the first row (chair class), the training curve of the baseline pseudo-label accuracy exhibits significant fluctuations, particularly showing a sudden sharp performance drop at approximately 50 epochs.
+Although the pseudo-label accuracy recovers slightly thereafter, the validation score does not.
+In the second row (sofa class), the training curve of the baseline exhibits more severe fluctuations and sharper and more drastic performance drops than in the first row.
+This phenomenon can be attributed to catastrophic forgetting where previously learned knowledge is forgotten when acquiring new knowledge [mccloskey1989catastrophic, kirkpatrick2017overcoming].
+In other words, self-training and class-imbalance scenarios (i.e., in our scenario) lead to instability in pseudo-label predictions, particularly for poorly behaved classes, due to the lack of labeled data.
+In contrast to the baseline, the proposed method shows a smoother training curve without significant fluctuations in either category.
+This indicates that the proposed training procedure aids in achieving stable optimization for poorly behaved classes that suffer from the forgetting problem (Refer to Sec.
+8 in the supplementary material for additional analysis).
+Figure 2: Training curves for the chair and sofa classes, illustrating variations in pseudo-label pixel accuracy and validation IoU scores (on Pascal VOC 92-label partition).
+Method Encoder Pascal VOC $\rightarrow$ COCO
+UniMatch [yang2023revisiting] RN-50 71.9 35.6
++ PrevMatch RN-50 73.4 (+1.5) 38.5 (+2.9)
+UniMatchV2-B [yang2023revisiting] DINOv2-B 86.3 52.9
++ PrevMatch DINOv2-B 87.2 (+0.9) 55.1 (+2.2)
+Table 7: Evaluation results on seen (Pascal VOC) and unseen (COCO) datasets.
+COCO contains 80 object classes, 20 of which correspond to the categories in Pascal VOC.
+The remaining 60 classes were treated as a background class for evaluation.
+The results on COCO were obtained without re-training, using the weights of the model trained on Pascal VOC with a 92-label setting.
+5.3 Training Stability and Generalization
+Several studies [li2018visualizing, foret2020sharpness, chen2021vision, shin2023metricgan] across various domains have demonstrated that stable optimization during training results in a flatter loss landscape, thereby improving generalization performance.
+Therefore, to assess whether the proposed method’s training stability contributes to improved generalization, we conduct additional experiments.
+Tab.
+7 presents evaluation results for models trained on Pascal VOC [everingham2010pascal] and tested on the COCO dataset [lin2014microsoft] without re-training.
+The results show a larger difference in the score on the unseen dataset (COCO) as compared to that on the seen dataset (Pascal VOC), indicating that PrevMatch improves the model’s generalization ability.
+Method #Trainable Networks Training Time (m) GPU Mem. Usage (G) Pascal
+92 183 366
+PS-MT 1 24.9 28.6 65.8 69.6 76.6
+Diverse Co-T. 3 22.1 40.5 75.4 76.8 79.6
+UniMatch 1 7.8 21.9 75.2 77.2 78.8
++ PrevMatch 1 9.1 22.7 77.0 78.5 79.6
+Table 8: Comparison of time spent using existing methods with a ResNet-101 encoder.
+Training time per epoch and GPU memory usage were measured using the same environment (two A6000 GPUs) and hyperparameters.
+We used the open-sourced code provided by the authors.
+5.4 Efficiency Evaluation
+To evaluate the efficiency of PrevMatch, we compare its training time per epoch and GPU memory usage with existing methods.
+Since UniMatch [yang2023revisiting] relies solely on input separation, it may suffer from instability due to the inherent coupling problem.
+In contrast, PS-MT [liu2022perturbed], Diverse Co-training [li2023diverse], and our proposed PrevMatch address this limitation by integrating both network and input separation techniques.
+In Tab.
+8, PrevMatch improves UniMatch by 1.8% (92-label setting), with a slight cost increase.
+Moreover, PrevMatch outperforms Diverse Co-training, while achieving 2.4 times faster training.
+Meanwhile, PS-MT, a dual Mean Teacher-based method, requires training time comparable to that of Diverse Co-training due to its complex components, despite being a single trainable network.
+These results suggest that PrevMatch can efficiently mitigate the coupling problem.
+Therefore, owing to its computational efficiency and the simplicity of its pipeline reusing the inputs and networks, the proposed method can be easily integrated into any semi-supervised learning method.
+6 Conclusion
+In this work, we introduced the PrevMatch framework, which leverages temporal knowledge obtained during training to efficiently address the issues of tight coupling and confirmation bias that impede stable semi-supervised learning.
+The main contributions of PrevMatch include revisiting the use of temporal knowledge and maximizing its effectiveness.
+Experiments were conducted on three benchmark datasets, revealing that incorporating PrevMatch into existing methods significantly improves their performance.
+In addition, our findings indicate that leveraging temporal knowledge facilitates stable optimization, particularly for classes that exhibit poor performance and fluctuations.
+Finally, the computational efficiency and compatibility of the proposed method facilitate its seamless integration into recent semi-supervised semantic segmentation methods.
+Acknowledgements.   This research was supported by the BK21 FOUR funded by the Ministry of Education of Korea and National Research Foundation of Korea.
+[…]
+Supplementary Material
+7 Additional Experiments and Ablation Studies
+In this study, experiments were conducted in the following environments: Ubuntu 20.04, Python 3.10.4, PyTorch 1.12.1, and NVIDIA 3090Ti or A6000 GPUs.
+Unless otherwise specified, subsequent experiments were primarily conducted using the UniMatch baseline with ResNet-50.
+7.1 Comparison with State-of-the-Art on the Priority Protocol of Pascal VOC
+As described in the main paper, the Pascal VOC dataset contains a high-quality annotation subset (1,464 images) and a coarse-quality annotation subset (9,118 images).
+In addition to the Original and Blended protocols presented in the main manuscript, the Priority protocol has been used in several studies for comparison.
+The protocol where the selection of labeled images is first derived from the high-quality subset; if not sufficient, it is complemented by additional images from the coarse subset.
+Tab.
+9 shows that the proposed method consistently improves baseline methods.
+7.2 Additional Ablation Studies
+Due to space constraints, the ablation study (Tab.
+5) in the main body of the paper includes only the 92-label protocol.
+To further validate the proposed method, we provide additional experiments on the 183- to 1464-label protocols.
+Tab.
+10 shows that the components of PrevMatch consistently improve performance across various evaluation protocols.
+7.3 Previous List Length
 We investigate the effect of the length ($N$) of the previous list that stores the temporal models.
-In Table VII, the results for $N=1$ and $N=2$ are comparable to those of the baseline.
+In Tab.
+11, the results for $N=1$ and $N=2$ are comparable to those of the baseline.
 This suggests that the aforementioned coupling problem may persist because the previous models in the list are continually updated with the latest model when $N$ is small.
@@
 However, the performance for $N=20$ increases only marginally due to the use of outdated teachers.
-Based on this result, we recommend setting the value of $N$ to approximately 5–15% of the total training epochs, which proves to be appropriate for different datasets (e.g., Pascal=6–10, Cityscapes=8–16, and COCO and ADE20k=4–5).
-TABLE VII: Ablation study for the maximum length of the previous list using a ResNet-50 encoder. In this setting, only previous guidance is used (i.e., $K$=1, without ensemble).
-List Length ($N$) Base. 1 2 4 8 12 20
+Based on this result, we recommend setting the value of $N$ to approximately 5–15% of the total training epochs, which proves to be appropriate for different datasets (e.g., Pascal=6–10, Cityscapes=8–16, and ADE20k=4–6).
+Pascal [Priority set] Encoder 1/16 1/8 1/4
+U2PL [wang2022semi] RN-101 77.2 79.0 79.3
+U2PL+ [wang2024using] RN-101 77.2 79.4 80.2
+Dual Teacher [na2023switching] RN-101 80.1 81.5 80.5
+CorrMatch [sun2024corrmatch] RN-101 81.3 81.9 80.9
+AllSpark [wang2024allspark] MiT-B5 81.6 82.0 80.9
+FixMatch [sohn2020fixmatch] RN-50 75.4 76.5 76.9
++ PrevMatch RN-50 76.9 77.6 77.8
+     Gain ($\Delta$) (+1.5) (+1.1) (+0.9)
+UniMatch [yang2023revisiting] RN-101 80.9 81.9 80.4
++ PrevMatch RN-101 81.4 81.9 80.8
+     Gain ($\Delta$) (+0.5) (+0.0) (+0.4)
+UniMatchV2-S [yang2025unimatch] DINOv2-S 86.6 87.3 85.9
++ PrevMatch DINOv2-S 87.3 87.8 86.4
+     Gain ($\Delta$) (+0.7) (+0.5) (+0.5)
+Table 9: Comparison with state-of-the-art methods on the Priority protocol of Pascal VOC dataset.
+Previous Random Random PASCAL
+Guidance Selection Weights 92 183 366 732 1464
+- - - 71.9 72.5 76.0 77.4 78.7
+- - 72.7 73.8 76.6 78.1 79.0
+- 73.2 74.9 77.4 78.3 79.2
+73.4 75.4 77.5 78.6 79.3
+Table 10: Additional ablation studies across different evaluation protocols (UniMatch with ResNet-50).
+NN Base. 1 2 4 8 12 20
 Pascal92 71.9 71.7 71.7 72.4 72.7 72.5 71.9
 Pascal183 72.5 72.7 72.9 73.4 73.8 73.5 73.3
-4.3.3 Upper Bound Number for Random Selection
+Table 11: Ablation study for the maximum length ($N$) of the previous list. In this setting, only previous guidance is used (i.e., KK=1, without ensemble).
+7.4 Upper Bound Number for Random Selection
 To generate reliable and diverse pseudo-labels, we proposed a strategy that randomly selects $k$ models (ranging from 1 to $K$) for each iteration.
 In this strategy, we explore the performance changes regarding the upper bound number $K$.
-Table VIII indicates that including the ensembling cases ($K>1$, i.e., $k$=1 or $k>$1 are randomly selected) improves the performance significantly compared to the case of $K=1$ (i.e., without ensemble).
+Tab.
+12 indicates that including the ensembling cases ($K>1$, i.e., kk=1 or k>k>1 are randomly selected) improves the performance significantly compared to the case of $K=1$ (i.e., without ensemble).
 In addition, we observe the best results at $K=3$ and a slight performance drop in settings with $K$ greater than 3.
 Even for large $K$, a varying number ($k$) of models is selected; however, the proportion of large $k$ values increases with $K$.
-This ensures consistent pseudo-labels but reduces their diversity, potentially degrading performance, as mentioned in Dual Teacher [31].
+This ensures consistent pseudo-labels but reduces their diversity, potentially degrading performance, as mentioned in Dual Teacher.
 In conclusion, we select $K=3$ because it adequately satisfies the diversity and reliability requirements of the pseudo-labels.
-TABLE VIII: Ablation study on the efficacy of the upper bound number $K$ using a ResNet-50 encoder and $N$=8.
 Upper Bound Number ($K$) 1 2 3 4 5
@@
 Pascal183 73.8 74.9 75.4 75.2 75.1
-4.3.4 Criteria for Saving Previous Models
-As described in Section 3.2, one alternative for storing previous models involves saving the model at regular intervals, a method used in [62, 63, 21].
+Table 12: Ablation study on the efficacy of the upper bound number $K$ using NN=8.
+Save Criteria 92 183 366 732 1464
+(a) Baseline 71.9 72.5 76.0 77.4 78.7
+(b) Every 1 Epoch 71.8 73.5 76.3 77.5 78.6
+(c) Every 3 Epochs 72.2 74.7 76.8 78.0 78.6
+(d) On Best Epochs (ours) 73.4 75.4 77.5 78.6 79.3
+Table 13: Ablation study regarding the efficacy of the save criteria using NN=8 and KK=3.
+7.5 Criteria for Saving Previous Models
+As described in Sec.
+3.2.1 of the main paper, one alternative for storing previous models involves saving the model at regular intervals, a method used in [huang2017snapshot, wang2022efficient, yang2022st++].
 Thus, we conduct experiments to validate the effectiveness of this approach.
-As listed in Table IX, although case (b) exhibits slightly better overall performance than the baseline (a), the difference is marginal.
+As listed in Tab.
+13, although case (b) exhibits slightly better overall performance than the baseline (a), the difference is marginal.
 This suggests that storing models at short intervals does not address the coupling problem between the teacher and student networks.
@@
 In addition, our approach does not require additional hyperparameter searches to determine appropriate intervals, thereby reducing unnecessary training costs.
-TABLE IX: Ablation study regarding the efficacy of the save criteria using a ResNet-50 encoder, $N$=8, and $K$=3.
-Save Criteria 92 183 366 732 1464
+7.6 Loss Weight
+The loss weight $\lambda$ for previous guidance plays a critical role in the overall training process.
+Since the primary goal of previous guidance is to mitigate confirmation bias–which tends to occur during the early to middle stages of training–its effect is most beneficial before the model becomes overconfident in incorrect predictions.
+However, in the early phase, the model’s predictions are typically unreliable.
+To address this, we employ a warmup schedule that gradually increases $\lambda$, similar to commonly used learning rate warmup strategies.
+Furthermore, as training progresses and the model becomes more stable and accurate, the need for strong regularization from previous guidance naturally decreases.
+To reflect this, we apply a decay to $\lambda$ in the later stages.
+Our final schedule (row (e)) allows previous guidance to act most strongly when it is most needed–during the middle of training–while reducing its influence as the model converges.
+Tab.
+14 presents an ablation study comparing different $\lambda$ scheduling strategies.
+The fixed (b) and linear decay (c) settings maintain a high weight even in the early stage, where the model’s predictions are still unstable.
+As a result, these configurations show limited improvements, likely due to the guidance based on noisy predictions.
+In contrast, the linear increase (d) and warmup+decay (e) strategies both incorporate a warmup phase, mitigating this issue and yielding significant performance gains across various label partitions.
+While both (d) and (e) outperform other variants, the warmup+decay strategy shows slightly better or comparable results, especially in high-label settings.
+This indicates that reducing the influence of previous guidance in the late stage, when the model is already well-trained, is beneficial for final performance.
+Loss weight ($\lambda$) 92 183 366 732 1464
 (a) Baseline 71.9 72.5 76.0 77.4 78.7
-(b) Every 1 Epoch 71.8 73.5 76.3 77.5 78.6
-(c) Every 3 Epochs 72.2 74.7 76.8 78.0 78.6
-(d) On Best Epochs (PrevMatch) 73.4 75.4 77.5 78.6 79.3
-Figure 3: Qualitative segmentation results on (a) Pascal VOC with a 92-label partition using a ResNet-50 encoder and (b) Cityscapes with a 1/16 label partition using a ResNet-101 encoder.
-4.4 Discussion
-TABLE X: Class-wise IoU scores for Pascal VOC with a 92-label partition using a ResNet-50 encoder. $\Delta$ indicates the difference between the PrevMatch and baseline UniMatch performances.
+(b) Fixed (1.0) 72.1 74.2 76.7 78.2 78.6
+[…]
+[…]
+[…]
+Table 14: Ablation study for $\lambda$.
+8 Additional Analysis
+Figure 3: Qualitative segmentation results on (a) Pascal VOC and (b) Cityscapes.
 backgr.
@@
 cow
-table
+d.table
 dog
@@
 person
-plant
+p.plant
 sheep
@@
 UniMatch 91 84 59 89 71 68 92 80 88 8 88 55 85 85 75 80 54 82 33 80 64
-PrevMatch 93 84 61 87 71 68 93 84 89 21 89 57 84 86 77 82 51 82 46 84 58
-$\Delta$ 2 0 2 -2 0 0 1 4 1 13 1 2 -1 1 2 2 -3 0 13 4 -6
-TABLE XI: Class-wise IoU scores for Cityscapes with a 1/8 label partition using a ResNet-101 encoder.
++ PrevMatch 93 84 61 87 71 68 93 84 89 21 89 57 84 86 77 82 51 82 46 84 58
+[…]
+Table 15: Class-wise IoU scores for Pascal VOC using a ResNet-50 encoder.
 road
@@
 UniMatch 98 82 92 56 60 62 71 79 92 60 95 82 63 95 83 87 79 68 77
-PrevMatch 98 84 92 60 63 63 71 80 92 63 95 82 64 95 83 88 81 69 77
-$\Delta$ 0 2 0 4 3 1 0 1 0 3 0 0 1 0 0 1 2 1 0
-4.4.1 Analysis of Performance and Training Stability
-Class-wise IoU Scores and Qualitative Evaluation.
-Tables X and XI list the category-wise IoU scores.
-In particular, Table X on Pascal VOC shows that the proposed method achieves notable performance gains for the chair and sofa classes, which were particularly challenging for the UniMatch baseline.
-In addition, Table XI on the Cityscapes dataset shows that the proposed method achieves the largest performance improvements for the wall, fence, and terrain classes, which are the lowest performing among the 19 classes in UniMatch.
++ PrevMatch 98 84 92 60 63 63 71 80 92 63 95 82 64 95 83 88 81 69 77
+[…]
+Table 16: Class-wise IoU scores for Cityscapes using a ResNet-101 encoder.
+8.1 Class-wise IoU Scores and Qualitative Evaluation
+Tabs.
+15 and 16 list all category-wise IoU scores.
+In particular, Tab.
+15 on Pascal VOC shows that the proposed method achieves notable performance gains for the chair and sofa classes, which were particularly challenging for the UniMatch baseline.
+In addition, Tab.
+16 on the Cityscapes dataset shows that the proposed method achieves the largest performance improvements for the wall, fence, and terrain classes, which are the lowest performing among the 19 classes in UniMatch.
 In addition to the quantitative results, the qualitative results shown in Fig.
 3 corroborate these findings, revealing consistent improvements in the same categories.
-Thus, these results suggest that utilizing previous knowledge helps prevent the catastrophic forgetting problem described in Section 3.2, even in semi-supervised semantic segmentation scenarios.
-We further investigate this problem in the subsequent analysis.
-Training Stability of Poorly Behaved Classes.
-To further explore the catastrophic forgetting problem in poorly behaved classes, we visualize changes in terms of pseudo-label accuracy and validation IoU scores during training, as depicted in Fig.
-4.
-In the first row (chair class), the training curve of the baseline pseudo-label accuracy exhibits significant fluctuations, particularly showing a sudden sharp performance drop at approximately 50 epochs.
-Although the pseudo-label accuracy recovers slightly thereafter, the validation score does not.
-In the second row (sofa class), the training curve of the baseline exhibits more severe fluctuations and sharper and more drastic performance drops than in the first row.
-In contrast, the proposed method shows a smoother training curve without significant fluctuations in either category.
-This indicates that the proposed training procedure aids in achieving stable optimization for poorly behaved classes that suffer from the forgetting problem.
-Figure 4: Training curves for the chair and sofa classes, illustrating variations in pseudo-label pixel accuracy and validation IoU scores.
-The experiment is conducted on the 92-label partition of Pascal VOC.
-Training Stability Across Different Label Partitions.
+Thus, these results suggest that utilizing previous knowledge helps prevent the catastrophic forgetting problem, even in semi-supervised semantic segmentation scenarios.
+Figure 4: Training curves for different label partition settings on Pascal VOC.
+The X- and Y-axes represent epochs and validation mIoU, respectively.
+The square symbol ($\blacksquare$) denotes the epoch with the best performance.
+8.2 Training Stability Across Different Label Partitions
 Fig.
 4 illustrates the effect of the proposed method on the changes in the validation scores throughout the training process.
-In fewer label settings (92 and 183), the baseline (blue) exhibits significant fluctuations in terms of performance compared to the proposed method (orange).
+In fewer label settings (92 and 183), the UniMatch baseline (blue) exhibits significant fluctuations in terms of performance compared to the proposed method (orange).
 Moreover, when considering the epoch that achieves the best performance, the baseline method struggles to converge consistently across epochs and tends to become trapped in local minima prematurely.
@@
 In contrast, the proposed method consistently converges across epochs without significant fluctuations.
-Finally, the consistent outperformance of our method over the baseline across almost all training epochs in label partitions suggests that the positive effects of previous guidance outweigh any negative effects.
-Note that the positive and negative effects refer to the cases (3) and (2), respectively, described in the last paragraph of Section 3.2.
-The results presented in this subsection suggest the following implications: (1) Training with fewer labels leads to instability in pseudo-label predictions, particularly for poorly behaved classes.
-(2) The proposed training procedure aids in stabilizing pseudo-label predictions while mitigating the forgetting problem in poorly behaved classes.
-(3) In other words, leveraging previous knowledge helps alleviate the issues of tight coupling and catastrophic forgetting, which can hinder stable learning in semi-supervised segmentation settings.
-Figure 5: Training curves for different label partition settings on Pascal VOC.
-The X- and Y-axes represent epochs and validation mIoU, respectively.
-The square symbol ($\blacksquare$) denotes the epoch with the best performance.
-TABLE XII: Comparison of time spent using existing methods with a ResNet-101 encoder.
-Training time per epoch and GPU memory usage were measured using the same environment (two A6000 GPUs) and hyperparameters, such as batch size, in a 92-label partition.
-We used the open-sourced code provided by the authors.
-Method #Trainable Networks Training Time (m) GPU Memory Usage (G) Pascal
-92 183 366
-UniMatch 1 7.8 21.9 75.2 77.2 78.8
-Diverse Co-T. 3 22.1 40.5 75.4 76.8 79.6
-PS-MT 1 24.9 28.6 65.8 69.6 76.6
-ours ($K$=1) 1 8.4 22.7 76.4 77.5 79.3
-ours ($K$=3) 1 9.1 22.7 77.0 78.5 79.6
-4.4.2 Efficiency Evaluation
-To investigate the complexities of the proposed method, we measure the training time per epoch and GPU memory usage.
-In Table XII, Diverse Co-training, based on the co-training approach, slightly outperforms UniMatch, which operates on a single trainable network, in terms of performance.
-However, Diverse Co-training requires approximately three times more training time and twice as much memory, resulting in limited scalability.
-Moreover, PS-MT, a dual Mean Teacher-based method, requires training time comparable to that of Diverse Co-training due to its complex components, despite being a single trainable network.
-In contrast, the proposed method significantly outperforms existing methods, with a slight cost increase.
-In particular, PrevMatch ($K$=3) outperforms UniMatch by 1.8% in terms of mIoU (92 label setting) while requiring only a slight increase in complexities in training time and GPU usage, suggesting that PrevMatch can efficiently provide diverse pseudo-labels.
-Therefore, owing to its computational efficiency and the simplicity of its pipeline reusing the same inputs, the proposed method can be easily integrated into any semi-supervised learning method.
-5 Conclusion and Future Work
-In this work, we introduced the PrevMatch framework, which leverages temporal knowledge obtained during training to efficiently address the issues of tight coupling and confirmation bias that impede stable semi-supervised learning.
-The main contributions of PrevMatch include revisiting the use of temporal knowledge and maximizing its effectiveness.
-Specifically, we directly utilized previous models to provide additional pseudo-label guidance, referred to as previous guidance, to the student network.
-In addition, we developed a highly randomized ensemble strategy that enhances the reliability and diversity of the previous guidance while minimizing the increase in computational complexity.
-Experiments were conducted on four benchmark semantic segmentation datasets, revealing that PrevMatch significantly outperforms existing state-of-the-art methods across different evaluation protocols.
-Furthermore, our findings indicate that leveraging temporal knowledge facilitates stable optimization, particularly for classes that exhibit poor performance and fluctuations.
-Finally, the computational efficiency and compatibility of the proposed method facilitate its seamless integration into recent semi-supervised semantic segmentation methods.
-In future work, we will investigate the capability of PrevMatch to address the domain adaptation problem.
-This problem poses a more challenging self-training task due to domain discrepancies between the labeled and unlabeled images, a commonly encountered problem in real-world applications.
+Meanwhile, as mentioned in Sec.
+3.2 of the main paper, given the standard and previous guidance, four scenarios can be considered based on whether they are correct or incorrect (standard-previous): (1) correct-correct, (2) correct-incorrect, (3) incorrect-correct, and (4) incorrect-incorrect.
+Through case (3), the network receives an additional opportunity to be guided in the right direction, away from the wrong one.
+However, model training may also be hindered through case (2), leading to significant fluctuations.
+Nevertheless, as shown in Fig.
+4, the consistent outperformance of our method over the baseline across almost all training epochs in different label partitions suggests that the positive effects (i.e., case (3)) of previous guidance outweigh any negative effects (i.e., case (2)).
+9 Limitations and Future Work
+We showed that the proposed method is effective on several benchmark datasets that share the domain between labeled and unlabeled images.
+However, it has not been investigated in domain adaptation problems that involve domain discrepancies between the labeled and unlabeled data.
+This problem poses a more challenging self-training task, a commonly encountered problem in real-world applications.
+In future work, we will investigate the capability of temporal knowledge in the domain adaptation problem.
 In addition, we observed a phenomenon where significant fluctuations in pseudo-label accuracy for several classes negatively affect generalization ability.
@@
 Therefore, we intend to explore this phenomenon extensively across different classes, in terms of the relationship between training instability and generalization ability.
-Acknowledgments
-This research was supported by Brain Korea 21 FOUR. This research was also supported by Korea University Grant (K2403371) and Korea TechnoComplex Foundation Grant (R2112653).